### v4

# Potato Potahto in the FAO-GAEZ Productivity Measures?

Abstract
The FAO-GAEZ productivity data are widely used in Economics. However, the empirical literature rarely discusses measurement error. We use two proxies to derive novel analytical bounds around the effect of agricultural productivity in a setting with nonclassical measurement error. These bounds rely on assumptions that are weaker than the ones imposed in empirical studies and exhaust the information contained in the first two moments of the data. We reevaluate three influential studies, documenting that measurement error matters and that the impact of agricultural productivity may be smaller than previously reported. Our methodology has broad applications in empirical research involving mismeasured variables.
Keywords: FAO-GAEZ; Crop Productivity; Measurement Error; Partial Identification.
JEL Codes: C13, C18, N50, O13, Q10
Introduction
The impact of agricultural productivity on economic outcomes is paramountly important in many fields, including Economic History (Nunn and Qian, 2011), Development Economics (Bustos et al., 2016), and Political Economy (Acharya et al., 2016). The most common measure of agricultural productivity is the Food and Agriculture Organization Global Agro-Ecological Zones (FAO-GAEZ) crop productivity measures. However, this measure is based on a prediction model and, as with any forecasting model, is subject to measurement error (Rhode, 2024). This type of complex challenge is acknowledged by FAO, which frequently updates its model and has done so four times (Fischer et al., 2021). However, the existence of measurement error in the FAO-GAEZ productivity data is rarely recognized in the empirical literature and is not properly addressed by current statistical methods.
To address this issue, we propose a novel analytical method that uses multiple proxies to partially identify the effect of agricultural productivity, accounting for measurement error. To illustrate the empirical relevance of our approach, we revisit three influential studies: Nunn and Qian (2011), Bustos et al. (2016), and Acharya et al. (2016). Applying our method with two available proxies, we find that the estimated effect of agricultural productivity on population might be substantially smaller than reported by Nunn and Qian (2011); that there might be no effect on political behavior, in contrast to the findings of Acharya et al. (2016); and that the effect on structural transformation remains robust, consistent with Bustos et al. (2016).
We start by documenting the possibility of nonclassical measurement error in the FAO-GAEZ productivity variables. The FAO-GAEZ project combines data on climate, soil, and plant growth parameters to predict potential productivity of various crops. However, these measures of agricultural productivity face challenges of model and data uncertainty. For example, FAO-GAEZ’s versions 3 and 4 differ with respect to calibrated parameters (e.g., the ability of different plants to capture sunlight) and sources of soil and climate data. These differences cause large changes in measured productivity for many crops and regions, suggesting the existence of nonclassical measurement error in the FAO-GAEZ data.
Motivated by our assessment of the empirical studies using the FAO-GAEZ productivity measures, we propose to partially identify the effect of agricultural productivity in a linear model with two proxies that are subject to nonclassical measurement error. In this linear model, we impose that the true unobservable productivity is exogenous and that its effect on the outcome is non-negative. [^1] Knowledge of the direction of the coefficient of interest may be motivated by economic theory or by assuming that measurement error issues are not severe enough to change the sign of the estimated coefficient in the empirical literature. If a researcher is not willing to impose this type of assumption, Appendix D proposes an alternative way to partially identify this coefficient without making assumptions about its sign. We also impose that the measurement errors in the two productivity proxies are exogenous in the sense that they are uncorrelated with the economic shock of the outcome variable. Importantly, we allow these measurement errors to be nonclassical and only restrict them to be positively correlated with each other. [^2] Positive correlation between the measurement errors in the two proxies is motivated by arguing that unobservable natural productivity factors in the FAO-GAEZ version 4 model are also unobservable in the FAO-GAEZ version 3 model (Appendix I). If a researcher is not willing to impose this type of assumption, Appendix C proposes an alternative way to partially identify this coefficient without restricting the correlation between the two measurement error terms.
Under these assumptions, we derive analytical bounds around the true linear effect of agricultural productivity on the outcome of interest. Importantly, our assumptions are no stronger than those commonly applied when ignoring measurement error, and our bounds exhaust all the information contained in the first two moments of the data distribution. Intuitively, our bounds create the shortest possible interval that contains the true target parameter and is constructed through restrictions on only the first two moments of the latent variables. Importantly, we focus on this type of restriction because most applied researchers use linear regression models and do not restrict higher-order moments.
Moreover, our bounds have intuitive closed-form solutions. First, the upper bound captures an extreme scenario when the proxy variables are uninformative about the true productivity, i.e., they contain only noise. Second, the lower bound collapses to the standard OLS estimand when there is no measurement error. As a consequence of this necessary condition, the similarity between these two objects may be used as a suggestive assessment of whether measurement error is a relevant issue in empirical applications.
Additionally, these bounds present two interesting features. First, they offer a test for the identifying assumptions. Second, they compare favorably against two alternative strategies in the literature: a reduced-form approach that ignores measurement error and a bias-minimizing tool recommended by Lubotsky and Wittenberg (2006). In particular, there exist data-generating processes where our bounds are closer to the true target parameter than the alternative tools.
We also propose estimation and inference methods for the identified bounds. The estimation method is parametric and easy to implement through linear regressions. The inference procedure adapts the tool proposed by Chernozhukov et al. (2013) and constructs confidence bounds around the identified set with a known confidence level.
To illustrate our methodology, we reevaluate the results of three empirical studies: Nunn and Qian (2011), Bustos et al. (2016) and Acharya et al. (2016). To do so, we replicate their findings using the reduced-form approach, which ignores measurement error, and compare these results with those obtained through our partial identification method. Although we find that the results by Bustos et al. (2016) are robust to nonclassical measurement error, our method reaches very different conclusions when compared to the findings of Nunn and Qian (2011) and Acharya et al. (2016). In particular, Nunn and Qian (2011) finds that the introduction of potatoes has a significant effect on population at the 1%-confidence level, and Acharya et al. (2016) finds that cotton suitability significantly increases Republican vote share at the 10%-confidence level. However, when we account for nonclassical measurement error, we do not reject the null of no effect at the confidence level used in each one of these applications.
Lastly, we highlight that our proposed methodology can be used in many empirical contexts. To employ our partial identification strategy, we only require an exogenous mismeasured treatment variable and two exogenous proxies. As examples of possible applications, we mention the impact of media access on economic outcomes (Olken, 2009; Adena et al., 2015) and the effect of subjective beliefs about life expectancy on economic behavior (Post and Hanewald, 2013; Gan et al., 2015; Bissonnette et al., 2017). In the first case, media access is measured through an irregular terrain model, which has been modified over time. In the second case, different imputation methods are used to address the focal answer problem when trying to measure subjective probabilities. Appendix F explains six other examples in the fields of Macroeconomics, Behavioral Economics, Economic History, and Political Economy.
Related Literature. This article contributes to two different branches of literature. Concerning its empirical contribution, many papers in economics have used the FAO-GAEZ data. Nunn and Qian (2011), in pioneering work, uses the potential productivity of potatoes to explain how its introduction in the Old World affected population and urbanization. Bustos et al. (2016); Bustos et al. (2019); Bustos et al. (2020) use the productivity of soybeans to measure the impact of introducing genetically modified seeds on structural transformation. Bentzen et al. (2017) uses productivity under rain-fed and irrigation conditions to study the impacts of irrigation on the origins of autocracies. Acharya et al. (2016) uses cotton productivity as an instrument for slavery to explain contemporary differences in political attitudes in the U.S. Carillo (2021) uses wheat productivity to study the impact of agricultural policies on industrialization in Italy. [^3] Similar papers include Moorthy (2022), Dias et al. (2023) and Carreira et al. (2024) on soybean productivity and gender disparity, water pollution, and deforestation; Galor and Özak (2016) on the productivity of crops available before and after the Columbian Exchange and time preferences; Alesina et al. (2011, 2013) on the productivity of various crops, plough use, and gender disparities; Dube et al. (2016) on maize productivity and the Mexican drug sector; Gollin et al. (2021) on the productivity of various crops and the impacts of the Green Revolution; Mayshar et al. (2022) on the relative productivity of cereals over roots and tubers and the origin of states; Fiszbein (2022) on the productivity of multiple crops and its effects on development. These papers use FAO-GAEZ data to estimate linear specifications of the impacts of crop productivity on various outcomes of interest. However, it is worth noting that there is a range of papers employing FAO-GAEZ data in different types of models, such as trade models (Costinot et al., 2016; Sotelo, 2020; Pellegrina, 2022; Farrokhi and Pellegrina, 2023), deforestation models (Sant’Anna, 2024; Araujo et al., 2025; Dominguez-Iino, 2021), and other types of spatial models (Adamopoulos and Restuccia, 2022; Adamopoulos et al., 2022; Imbert et al., 2022; Fernández-Villaverde et al., 2023).
We contribute to this literature by reevaluating the results of Nunn and Qian (2011), Bustos et al. (2016) and Acharya et al. (2016) while accounting for measurement error in the main explanatory variable. In particular, we find that the effect of agricultural productivity on economic outcomes might be smaller than previously reported.
Concerning its theoretical contribution, our work is inserted in the literature about identifying treatment effect parameters with measurement error. As illustrated by Schennach (2016) and Hu (2017), this literature is vast and growing. The closest paper to ours was written by Lubotsky and Wittenberg (2006). While we derive an identified set that contains the true target parameter, they propose a single estimand that minimizes bias. Moreover, they impose that the measurement error terms are uncorrelated with the true variable of interest, while we impose no restriction on this correlation. This extra flexibility is particularly important in our empirical application because measurement error terms must be correlated with the true variable of interest when the latter is bounded, as is the case with crop productivity.
We also highlight that our work is connected to the studies developed by Hausman et al. (1998), Black et al. (2000), Hu and Schennach (2008), Cunha et al. (2010), Schennach (2014), Hu and Sasaki (2015), Alix-García and Millimet (2023) and Millimet and Parmeter (2024). Black et al. (2000) and Millimet and Parmeter (2024) differ from our work because they impose that the true variable of interest and the measurement error must be negatively correlated with each other. Furthermore, Hausman et al. (1998) and Alix-García and Millimet (2023) focus on mismeasured binary outcome variables. Moreover, Hu and Schennach (2008), Cunha et al. (2010) and Hu and Sasaki (2015) require the measurement error terms to be mutually independent. Lastly, Schennach (2014) proposes a general method that encompasses our setting. However, her tool requires simulation-based optimization algorithms while our tool is based on an analytic solution that is easily implemented using standard statistical software. The importance of analytical solutions such as ours is emphasized by Schennach (2014, page 366).
Paper Organization. The remainder of this paper is organized as follows. Section 2 describes the FAO-GAEZ data, documents the existence of measurement error in its variables, and intuitively explains the use of partial identification as a solution to measurement error. Section 3 presents our econometric model and states our main identification results. Section 4 briefly describes the estimation and inference methods connected to our partial identification strategy. Section 5 illustrates the proposed tools by reevaluating the impact of agricultural productivity on three empirical contexts, and Section 6 explains how to use our tools in a practical way in other applications. Lastly, Section 7 concludes.
Measurement Error in the FAO-GAEZ Data and Partial Identification as a Solution
In this section, we explain the empirical context of our methodology and how it is connected with the proposed identification strategy. Section 2.1 explains the FAO-GAEZ data and the three empirical examples we discuss in this paper. Section 2.2 intuitively explains partial identification and its usefulness in the applied researcher’s toolkit.
Empirical Context and Data
The Food and Agriculture Organization Global Agro-Ecological Zones project (FAO-GAEZ) produces a widely used measure of crop productivity. To do so, it combines climate and soil data with plant growth parameters to simulate optimal growth cycles of crops to predict potential crop productivity globally at a resolution of approximately 10 km by 10 km (Fischer et al., 2012, 2021). The FAO-GAEZ generates productivity measures of crops under different scenarios of water source — rainfed or irrigation — and use of inputs — low use of inputs and high use of inputs, such as machinery, fertilizer, seeds and others. [^4] Before version 4, FAO-GAEZ contained a scenario with intermediate use of inputs. This option is no longer available in version 4.
There are two main advantages that led researchers to widely adopt the FAO-GAEZ data. First, it is independent of observed productivity, allowing it to capture potential productivity even in regions where specific crops are not cultivated. Second, it is not influenced by economic factors such as credit constraints, transportation infrastructure, and subsidies, which likely affect observed agricultural productivity. Thus, FAO-GAEZ is often interpreted as exogenous to institutional factors, greatly simplifying the interpretation of estimated parameters as being causal.
As the FAO-GAEZ data combine climate, soil, and crop growth parameters to build potential productivity measures, they are susceptible to nonclassical measurement error. Although this issue has been ignored in the literature, the FAO-GAEZ data offer a unique opportunity to better understand the structure of such measurement errors since the data has been updated and is currently on its fourth version.
There is no precise documentation of the differences between versions. Upon inspection of versions 3 and 4, we identified many differences in parameters and raw data. To illustrate this issue, we provide four examples. First, the “harvest index” parameter essentially measures the portion of the weight of the crop that is actually of economic interest — e.g., the beans of soybeans — and is different for a set of crops between versions. Second, for another set of crops, FAO-GAEZ updated the “leaf area index”, which measures the plant’s canopy and its ability to capture sunlight. Third, precipitation data in version 3 is from the Variability Analysis of Surface Climate Observations 1.1 (VASClimO), while version 4’s precipitation data is from the Global Precipitation Climatology Centre (GPCC) Full Data Reanalysis Product Version 6. Fourth, data that still uses the same source are using newer versions. For example, temperature, sunshine fraction, and wind speed were from CRU 1.1 in version 3 and from CRU 3.21 in version 4, which entails a range of internal modifications of each data source. [^5] Rhode (2024) discusses differences in the model for the productivity of cotton. Despite these differences, all versions of the FAO-GAEZ intend to measure the same potential productivity, i.e., the agricultural productivity for a given place, climate, technology, and water supply. Appendix I further discusses sources of measurement error in FAO-GAEZ data.
Rather than focusing on improving individual versions of the FAO-GAEZ data, our approach centers on combining two available proxies (versions 3 and 4) to identify the impact of the underlying agricultural productivity on economic outcomes. To do so, we use assumptions that are weaker than those commonly applied when using only one version. [^6] Version 1 was released in 2000, version 2 in 2002, version 3 in 2012, and version 4 in 2020. Although our framework can be extended to accommodate more than two proxies, we do not pursue this path for the FAO-GAEZ application because we do not have access to the data from versions 1 and 2.
To illustrate the importance of measurement error in FAO-GAEZ data and motivate our econometric framework in the next section, we highlight three applications in the literature: Nunn and Qian (2011), Bustos et al. (2016), Acharya et al. (2016). In the remainder of this section, we discuss some of the differences in agricultural productivity measured by different versions of FAO-GAEZ for the crops studied in these three articles. Importantly, the productivity measures used by Nunn and Qian (2011) and Acharya et al. (2016) are bounded, implying that any measurement error they may contain must be nonclassical.
Nunn and Qian (2011) study the impact of the introduction of the potato in the Old World on population, using the FAO-GAEZ data on suitability of white potato. Figure 1 shows the spatial differences between versions 3 and 4 for potato productivity. Some countries — such as Norway and Iraq — had little to no area suitable for potato in version 3 but a large suitable area in version 4, creating a large change in logs. Another example is China, which has an increase in area suitable for potatoes of 700% across versions.
Figure 1: Log of area suitable for potato in FAO-GAEZ versions 3 and 4
This figure shows the spatial distribution of the differences in FAO-GAEZ versions 3 and 4 for the log of the area suitable for potato (low input) on the left. The plot on the right shows the correlation between versions 3 and 4 for the log of area suitable for potato $(A^{potato})$. Notice we follow Nunn and Qian (2011) in using the transformation $\log(1+x)$ since our objective is that our results are comparable with the ones in the original paper. The 45 line is shown in red.
Bustos et al. (2016) studies the impact of the introduction of genetically modified soybean seeds in Brazil on structural transformation. To do so, the authors use the difference between soybean productivity under high and low inputs as a measure of productivity shock. Figure 2 shows the spatial differences between versions 3 and 4 colored by deciles of this productivity shock. The differences are highly heterogeneous in space: 10% of regions lose more than 120 kg of production per hectare while another 10% gain more than 500 kg per hectare from an average productivity before the shock of 300 kg per hectare in version 3.
Figure 2: Soybean productivity in FAO-GAEZ versions 3 and 4
This figure shows the spatial distribution of the differences in versions 3 and 4 for the change in soybean productivity (high input minus low input) on the left. Each color represents a decile of the data distribution. The plot on the right shows the correlation between versions 3 and 4 for the change in soybean productivity $(\Delta A^{soy})$. The 45 line is shown in red.
Lastly, Acharya et al. (2016) studies the impact of slavery on contemporary political attitudes in the South of the United States, using cotton productivity as an instrument for slavery. Figure 3 shows the spatial differences between versions 3 and 4 colored by quintiles of cotton productivity with high input. The differences are highly heterogeneous in space: 10% of regions lose more than 180 kg of production per hectare, which is close to 10% of the average productivity in version 3.
Figure 3: Cotton productivity in FAO-GAEZ versions 3 and 4
This figure shows the spatial distribution of the differences in versions 3 and 4 for the cotton productivity high input on the left. Each color represents a quantile of the data distribution. The plot on the right shows the correlation between versions 3 and 4 for cotton productivity $(A^{cotton})$. The 45 line is shown in red.
In summary, there is significant variation in the measured productivity of crops across FAO-GAEZ versions 3 and 4. These differences suggest the existence of nonclassical measurement error in the FAO-GAEZ productivity variables and motivate our econometric framework.
What is Partial Identification? Why is it useful?
To address the measurement error issue explained in the last section, we propose a novel partial identification strategy. Since this approach is not widely used in empirical studies, this section briefly discusses the concept of partial identification while Appendix G discusses it in more detail. The reader who is familiar with partial identification may skip this section entirely.
Empirical research relies on assumptions about a data-generating process, which models the relationship between latent and observable variables. Point identification imposes assumptions on this model to express the target parameter as a function of observable variables. This approach often requires strong restrictions that researchers may diverge about (Manski, 2011). Partial identification addresses this type of model uncertainty. A researcher may adopt a weaker set of assumptions, which is not sufficient for point-identification but is sufficient for partial identification. In this case, the researcher builds an interval that contains the target parameter and is expressed in terms of observables.
Different assumptions yield different interval lengths. In our context, one researcher may follow the literature and assume no measurement error; another (Section 3) is not willing to ignore measurement error but is willing to impose restrictions on the signs of the true effect and of the correlation between the measurement error terms; a third (Appendices C and D) is not willing to make one of these assumptions. The identified set of the first researcher is a point, of the second one is an interval, and of the third one is a wider interval. Importantly, the length of the interval is a result of model uncertainty. Sample uncertainty is accounted for by building a confidence set around the identified interval, similar to the point-identification approach that builds confidence intervals around the identified point.
Econometric Framework and Identification
Most papers using the GAEZ productivity measures (Nunn and Qian, 2011; Bustos et al., 2016; Acharya et al., 2016) are interested in a linear model similar to the following one:
$Y=\alpha_{0}+\sum_{j=1}^{J}\alpha_{j}\cdot W_{j}+\beta\cdot X^{*}+\epsilon,$ (1)
where $J\in\mathbb{N}$, $\alpha_{0},\ldots,\alpha_{J}$ are population parameters, $\beta$ is the population parameter of interest, $Y$ is the outcome variable (e.g., the change over time of the share of workers employed in manufacture in Bustos et al. (2016)), $W_{1},\ldots,W_{J}$ are the covariates (e.g., a time trend and municipality characteristics in Bustos et al. (2016)), $X^{*}$ is the true unobserved variable of interest (e.g., the true change in potential yield of soy after the introduction of new seeds in Bustos et al. (2016)) and $\epsilon$ is an unobservable economic shock that satisfies $Cov\left(A,\epsilon\right)=0$ for any $A\in\left\{W_{1},\ldots,W_{J},X^{*}\right\}$.
For brevity, we apply the Frish-Waugh-Lovell Theorem to rewrite Equation (1) as
$Y_{res}=\beta\cdot X_{res}^{*}+\epsilon_{res},$ (2)
where $Y_{res}\coloneqq Y-W\cdot\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot Y\right]$, $X_{res}^{*}\coloneqq X^{*}-W\cdot\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot X^{*}\right]$, $\epsilon_{res}\coloneqq\epsilon-W\cdot\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot\epsilon\right]=\epsilon$, and $W=\left[1,W_{1},\ldots,W_{J}\right]$ is the row vector of covariates including the constant term.
Since $X^{*}$ is unobservable, we cannot identify $\beta$ based on Equation (2) alone. To circumvent this issue, empirical researchers have employed a proxy variable: $Z_{1}$, which is usually constructed from the FAO-GAEZ data version 3 and is described in Section 2.1. One important point is that this proxy variable has been refined over time, resulting in the proxy constructed from the FAO-GAEZ data version 4. For this reason, we consider the use of two proxies — $Z_{1}$ and $Z_{2}$, representing FAO-GAEZ data versions 3 and 4 — that satisfy
$\begin{matrix}Z_{1,res}=X_{res}^{*}+U_{1}\\ Z_{2,res}=X_{res}^{*}+U_{2}\end{matrix},$ (3)
where $Z_{1,res}=Z_{1}-W\cdot\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot Z_{1}\right]$ and $Z_{2,res}=Z_{2}-W\cdot\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot Z_{2}\right]$. Intuitively, $Z_{1,res}$ and $Z_{2,res}$ are two imperfect measures of $X_{res}^{*}$ where $U_{1}$ and $U_{2}$ capture the measurement error.
Our goal is to use these two proxies to partially identify $\beta$. Section 3.1 states and intuitively explains our modeling assumptions, while Section 3.2 presents our main identification result.
Assumptions
To partially identify $\beta$, we impose six assumptions, contextualizing them to our empirical setting. For brevity, we impose assumptions directly on the residualized variables in Equations (2) and (3). In Appendix H, we impose assumptions on the primitives of our model (Equation (1)) that imply the assumptions described in this section.
Assumption 1 (Exogeneity)
After partialling out the covariate vector $W$, the true unobserved measure $X^{*}$ is exogenous, i.e., $Cov\left(X_{res}^{*},\epsilon_{res}\right)=0$.
Assumption 1 is a standard exogeneity condition in any linear model. For example, it is implied by the assumption that $Cov\left(A,\epsilon\right)=0$ for any $A\in\left\{W_{1},\ldots,W_{J},X^{*}\right\}$ and is imposed in the empirical studies discussed in this article (Nunn and Qian, 2011; Bustos et al., 2016; Acharya et al., 2016). In our empirical application, it intuitively imposes that the natural forces that determine crop productivity (e.g., climate and soil nutrients captured by $X_{res}^{*}$) are uncorrelated with the unobserved economic forces that determine the outcome variable (e.g., financial shocks captured by $\epsilon_{res}$) after controlling for observable economic forces (e.g., educational levels captured by $W$).
Assumption 2 (Scale Normalization)
The scale of the true unobserved measure is normalized to have unit variance, i.e., $Var\left(X_{res}^{*}\right)=1$.
Assumption 2 allows us to interpret $\beta$ as the effect of increasing one standard deviation of $X_{res}^{*}$ on the outcome variable. Since $X_{res}^{*}$ is unobservable, this restriction is a simple normalization: if the true model were
$Y_{res}=\tilde{\beta}\cdot\tilde{X}_{res}^{*}+\epsilon_{res}$
with $Var\left(\tilde{X}_{res}^{*}\right)\neq 1$, we could rewrite it according to Equation (2) by defining
$\beta\coloneqq\tilde{\beta}\cdot\sqrt{Var\left(\tilde{X}_{res}^{*}\right)}\text{ and }X_{res}^{*}\coloneqq\dfrac{\tilde{X}_{res}^{*}}{\sqrt{Var\left(\tilde{X}_{res}^{*}\right)}},$
enforcing Assumption 2. One may view this normalization as changing the target parameter from $\tilde{\beta}$ to $\beta=\tilde{\beta}\cdot\sqrt{Var\left(\tilde{X}_{res}^{*}\right)}$ since it is not possible to separately identify $\tilde{\beta}$ and $\sqrt{Var\left(\tilde{X}_{res}^{*}\right)}$.
Assumption 3 (Known Direction of the True Effect)
The direction of the effect of the true unobservable variable of interest on the outcome variable is known, i.e., $\beta\geq 0$.
Assumption 3 imposes that the effect of the true variable of interest on the outcome variable is non-negative. [^7] The case that imposes that the true target parameter is non-positive is symmetric. Importantly, this assumption does not rule out that the effect might be zero. In our empirical application, it imposes that the effect of the true crop productivity on the outcome variable is non-negative. This assumption is motivated by the empirical literature because we assume that the articles using one proxy — the FAO-GAEZ version 3 — are at least identifying the correct sign of the true effect. If a researcher is not willing to impose this type of assumption, Appendix D proposes an alternative way to partially identify this coefficient without making assumptions about its sign.
Assumption 4 (Exogeneity of the Measurement Errors)
The measurement error variables are exogenous, i.e., $Cov\left(U_{1},\epsilon_{res}\right)=0$ and $Cov\left(U_{2},\epsilon_{res}\right)=0$.
Assumption 4 imposes that the measurement error terms in our proxies of the true variable of interest are uncorrelated with the unobservable economic shock after controlling for observable covariates. [^8] This assumptions is implied by $Cov\left(A,\epsilon\right)=0$ for any $A\in\left\{W_{1},\ldots,W_{J},X^{*},Z_{1},Z_{2}\right\}$ as discussed in Appendix H. This assumption weakens the implicit identifying assumption in the empirical literature about agricultural productivity (Nunn and Qian, 2011; Bustos et al., 2016; Acharya et al., 2016) since applied researchers impose that $U_{1}=U_{2}=0$.
In our empirical application, Assumption 4 intuitively imposes that prediction errors in the FAO-GAEZ measures are uncorrelated with the unobserved economic forces that determine the outcome variable. This assumption would be violated if the institutional factors that impact $\epsilon_{res}$ positively also affect the quality of the environmental data that is used when constructing the FAO-GAEZ measure. In this case, regions with large values of $\epsilon_{res}$ would have a small amount of measurement error. Consequently, $Cov\left(U_{1},\epsilon_{res}\right)=0$ would be implausible in this hypothetical scenario.
Assumption 5 (Nonclassical Measurement Error)
The measurement error terms are positively correlated with each other, i.e., $Cov\left(U_{1},U_{2}\right)\geq 0$. We name this assumption nonclassical measurement error because the covariances between the true variable of interest and the measurement error terms are unrestricted, i.e., $Cov\left(X_{res}^{*},U_{k}\right)\in\mathbb{R}$ for any $k\in\left\{1,2\right\}$.
Assumption 5 imposes that the covariance between the measurement error terms is positive. In our empirical application, it intuitively imposes that areas with large prediction errors in the FAO-GAEZ version 3 model tend to have large prediction errors in the FAO-GAEZ version 4 model too. This assumption is plausible because unobservable natural productivity factors in the FAO-GAEZ version 4 model are also unobservable in the FAO-GAEZ version 3 model. [^9] Appendix C offers an alternative partial identification strategy that does not restrict the covariance between the two measurement error terms. As a consequence of imposing fewer assumptions, the identified interval around the parameter of interest is weakly wider. In Appendix I, we discuss sufficient conditions in the FAO-GAEZ data under which this assumption holds.
Moreover, Assumption 5 allows nonclassical measurement error. Importantly, it does not restrict the relationship between the true variable of interest and the measurement error terms. Such flexibility is empirically relevant in many possible scenarios. For example, our framework allows the models used to construct the proxies to underpredict areas with high productivity and overpredict areas with low productivity. In this scenario, there is nonclassical measurement error because $Cov\left(X_{res}^{*},U_{k}\right)<0$ for any $k\in\left\{1,2\right\}$.
Assumption 6 (Well-behaved Data Distributions)
The variance of the outcome variable is strictly positive, the covariances between each proxy and the outcome variable are positive, and the covariance of the two proxy variables is strictly greater than the negative of the variable of interest’s variance, i.e., $Var\left(Y_{res}\right)>0$, $Cov\left(Z_{1,res},Y_{res}\right)\geq 0$, $Cov\left(Z_{2,res},Y_{res}\right)\geq 0$ and $Cov\left(Z_{1,res},Z_{2,res}\right)>-1.$
Assumption 6 imposes four restrictions on the distribution of the observable variables. The first restriction imposes that the outcome variable is not a constant. If we do not impose this assumption, we must have that $\beta=0$ by construction and the identification problem in Equation (2) is uninteresting. The second and third conditions impose that the reduced form effects have the same sign as the true target parameter, while the fourth condition imposes that the correlation between the two proxies is sufficiently large. These restrictions are plausible in our empirical application because both proxies are trying to predict the same underlying true productivity. Furthermore, note that all restrictions in Assumption 6 are easily testable.
Identification
In this section, we state our main identification result. Before doing so, we define two objects
$LB\coloneqq\max\left\{\begin{matrix}\dfrac{Cov\left(Z_{1,res},Y_{res}\right)+Cov\left(Z_{2,res},Y_{res}\right)}{1+Cov\left(Z_{1,res},Z_{2,res}\right)},\\ \\ 2\cdot\dfrac{Cov\left(Z_{1,res},Y_{res}\right)}{1+Var\left(Z_{1,res}\right)},\hskip 10.0pt2\cdot\dfrac{Cov\left(Z_{2,res},Y_{res}\right)}{1+Var\left(Z_{2,res}\right)}\end{matrix}\right\}$ (4)
and
$UB\coloneqq\sqrt{Var\left(Y_{res}\right)},$ (5)
and note that $LB\geq 0$ according to Assumption 6.
We now state our main result: partial identification of $\beta$ (Equation (2)) using two proxies (Equation (3)) with nonclassical measurement error.101010 As discussed in Section 2.2, the bounds in Proposition 1 account for model uncertainty in the sense that the researcher is not sure about the magnitude of the correlation between the true unobservable variable of interest and the proxies’ measurement error terms. Due to the uncertainty behind the underlying model, the researcher may prefer to use the bounds in Proposition 1 or the ones in Appendices C and D, which use different set of assumptions. Importantly, these bounds treat the populational distribution of the observable variables as known objects. Consequently, they do not account for sample uncertainty. This type of uncertainty will be addressed in Section 4.
Proposition 1
Under Assumptions 1-6, we have that:
$\beta\in\left\{\begin{array}[]{cl}\left\{\emptyset\right\}&\text{ if }UB<LB\\ \left[LB,UB\right]&\text{ if }UB\geq LB\end{array}\right.$. (6)
Moreover, the bounds in Equation (6) exhaust all the information contained in the first two moments of the observable variables $\left(Y_{res},Z_{1,res},Z_{2,res}\right)$. [^11] See Appendix A.1.4 for the formal definition of “exhausting all the information contained in the first two moments of the observable variables.” Below, we provide a intuitive definition for this concept.
Proof.
See Appendix A.1. ∎
Proposition 1 has two intuitive aspects. One is connected to the lower bound while the other one is connected to the upper bound.
First, the lower bound in Equation (4) collapses to the standard OLS estimand when there is no measurement error. Consequently, similarity between the estimated lower bound and the estimated OLS coefficient may be used as an indirect test of whether measurement error is a relevant issue in a given empirical application. Importantly, this test relies on a necessary condition of “no measurement error” instead of a sufficient condition. For this reason, it should be used as a suggestive assessment only, and it should be combined with an economic argument that explains why measurement error is irrelevant in a specific application.
Second, the upper bound captures an extreme scenario for measurement error. It is connected to the case when the proxy variables are composed of noise only. The upper bound reflects this uninformative scenario because it does not depend on the proxy variables.
Furthermore, Proposition 1 has three important consequences that need further discussion: (i) it offers a test for our identifying assumptions’ validity, (ii) it exhausts all the information contained in the first two moments of the data distribution, and (iii) it compares favorably against alternative strategies in the literature.
First, if $UB<LB$, Proposition 1 implies that there is no data-generating process that satisfies Assumptions 1-6 and the restrictions imposed by the first and second moments of the observable data distribution. Consequently, $UB\geq LB$ is a testable necessary condition for the validity of our identifying assumptions.
Second, the bounds in Equation (6) exploit all the restrictions imposed by $\mathbb{E}\left[Y_{res}\right]$, $\mathbb{E}\left[Z_{1,res}\right]$, $\mathbb{E}\left[Z_{1,res}\right]$, $Var\left(Y_{res}\right)$, $Var\left(Z_{1,res}\right)$, $Var\left(Z_{2,res}\right)$, $Cov\left(Y_{res},Z_{1,res}\right)$, $Cov\left(Y_{res},Z_{2,res}\right)$, $Cov\left(Z_{1,res},Z_{2,res}\right)$ on the data generating process. Since they do not use restrictions imposed by all the moments of the data distribution, they are not sharp as defined by Canay and Shaikh (2017). However, in a linear model context, most researchers are willing to impose restrictions, at most, on the second moments of the underlying data-generating process, and our identification strategy exhausts the information contained in these types of moments. [^12] A similar approach was adopted by Mogstad et al. (2018) in a different setting. Focusing on an IV problem, their partial identification results (Mogstad et al., 2018, Proposition 3) exploit all the information contained in the conditional expectation of the outcome variable given the instrument, but do not exploit higher-order moments.
Third, our method compares favorably against two alternative strategies in the literature. The first alternative method is a reduced-form approach that ignores measurement error. For example, suppose a researcher performs the following regression:
$Y_{res}=b_{0}+b_{1}\cdot Z_{1,res}+V_{1},$ (7)
where $Cov\left(Z_{res},V_{1}\right)=0$ by construction. The second alternative method is recommended by Lubotsky and Wittenberg (2006) and tries to use convex combinations of the proxy variables to find an OLS coefficient as close as possible to the target parameter. In Appendix B, we use three simple data-generating processes to show that (i) $b_{1}$ may be strictly smaller than our lower bound, (ii) $b_{1}$ may be strictly larger than our upper bound, and (iii) our lower bound may achieve a smaller bias than the procedure proposed by Lubotsky and Wittenberg (2006).
Remark 1
If a researcher is willing to impose more assumptions on the data-generating process, she may tighten the bounds derived in Proposition 1. For example, imagine that the second proxy, $Z_{2,res}$, is better than the first proxy, $Z_{1,res}$, in the sense that $Var\left(U_{1}\right)\geq Var\left(U_{2}\right)$. Under this additional assumption, we tighten the bounds in Equation (6) with the following restriction:
$\left\{\begin{array}[]{rclcl}\beta&\geq&2\cdot\left[\frac{Cov\left(Z_{1,res},Y_{res}\right)-Cov\left(Z_{2,res},Y_{res}\right)}{Var\left(Z_{1,res}\right)-Var\left(Z_{2,res}\right)}\right]&\text{if}&Var\left(Z_{1,res}\right)>Var\left(Z_{2,res}\right)\\ \beta&\leq&2\cdot\left[\frac{Cov\left(Z_{1,res},Y_{res}\right)-Cov\left(Z_{2,res},Y_{res}\right)}{Var\left(Z_{1,res}\right)-Var\left(Z_{2,res}\right)}\right]&\text{if}&Var\left(Z_{1,res}\right)<Var\left(Z_{2,res}\right)\end{array}\right.$.
Estimation and Inference
In this section, we briefly describe a simple parametric estimator for the bounds described in Proposition 1 and a method to compute $\alpha$-confidence sets that contain the identified set with a probability at least as larger as the chosen confidence level $\alpha\in\left(0,1\right)$. To do so, we assume that we observe an independently and identically distributed sample $\left\{Y_{i},W_{1,i},\ldots,W_{J,i},Z_{1,i},Z_{2,i}\right\}_{i=1}^{N}$, where $N\in\mathbb{N}$ is the sample size.
To estimate the bounds described in Proposition 1, we recommend taking the sample analogs of Equations (4) and (5). To do so, residualize the outcome and proxy variables by regressing them on the covariates. Then, use these residuals to compute the desired sample covariances and sample variances. See Appendix E.1 for a detailed explanation of our estimation procedure.
To construct an $\alpha$-confidence interval around the identified set, we combine the confidence bounds proposed by Chernozhukov et al. (2013) with a Bonferroni-style correction. Since the upper bound does not rely on intersection bounds, we use a simple bootstrap procedure to compute a $\frac{\left(1+\alpha\right)}{2}$-confidence bound above it. Since the lower bound relies on intersection bounds, we use the method proposed by Chernozhukov et al. (2013) to compute a $\frac{\left(1+\alpha\right)}{2}$-confidence bound below it. By taking the intersection of these two confidence bounds, we construct an $\alpha$-confidence interval around the identified set. See Appendix E.2 for a detailed explanation of our inference procedure.
Empirical Results
To illustrate the use of our partial identification strategy, we reevaluate the findings of Nunn and Qian (2011), Bustos et al. (2016), and Acharya et al. (2016). We find that the estimated effect of agricultural productivity on population may be substantially smaller than reported by Nunn and Qian (2011); that there may be no effect on political behavior, in contrast to the findings of Acharya et al. (2016); and that the effect on structural transformation remains robust, consistent with Bustos et al. (2016).
To map our assumptions to the three papers, we note the following. Assumption 1 (Exogeneity) is imposed in all of them. Assumption 2 (Scale Normalization) only affects the interpretation of results, which are expressed in terms of standard deviations of $X^{*}$. Assumption 3 (Known Direction of the True Effect) requires that each application correctly identifies at least the sign of the effect. [^13] Appendix D shows how this assumption can be relaxed in our setting. Assumption 4 (Exogeneity of Measurement Errors) weakens the identification requirements relative to those in the papers. Assumption 5 (Nonclassical Measurement Error) is motivated by the persistence of measurement error in FAO-GAEZ data and is further discussed in Appendix I. Finally, Assumption 6 (Well-behaved Data Distributions) is a standard technical condition.
For other possible applications using different datasets, see Appendix F.
Nunn and Qian (2011)
Nunn and Qian (2011) study the impact of the introduction of the potato in the Old World on population. The empirical specification uses panel data to estimate Equation (8):
$y_{it}=\alpha_{i}+\alpha_{t}+\beta\cdot I^{Post}\cdot\log(1+A_{i}^{potato})+\sum_{j=1}^{J}\alpha_{j}\cdot w_{i,j}+\epsilon_{it},$ (8)
where $y_{it}$ is population in country $i$ and year $t$, $A_{i}^{potato}$ is the area of country $i$ that is suitable for the cultivation of white potato, [^14] A pixel is defined as being suitable if it has $\geq 40\%$ of the maximum attainable yield for the crop. $I^{Post}$ is a dummy variable that equals one for years after 1700, i.e., after the introduction of the potato. In our setting, the interaction $I^{Post}\cdot\log(1+A_{i}^{potato})$ is the variable of interest $X^{*}$.
A practical problem in replicating the work of Nunn and Qian (2011) is that the original results of Nunn and Qian (2011) use the intermediate input setting of FAO-GAEZ version 2. The FAO-GAEZ version 2 is not available anymore and the intermediate input specification is not available for version 4. For this reason, we replicate the results using the low input setting of versions 3 and 4 to guarantee a better comparison across versions.
Table 1 shows descriptive statistics for the interaction between the Post-1700 indicator variable and the natural logarithm of the area suitable for potatoes as measured by the FAO-GAEZ proxies. In version 4, the mean, standard deviation and maximum of this interaction increase. Consequently, there is more spatial heterogeneity in version 4 than in version 3.
Table 1: Descriptive Statistics: Area Suitable for Potato (Interaction)
$I^{Post}\cdot\log(1+A_{i}^{potato})$ $\Delta\left\{I^{Post}\cdot\log(1+A_{i}^{potato})\right\}$
V 3 V 4 V 4 - V 3
(1) (2) (3)
Mean 1.31 1.56 0.25
S.D. 2.92 3.27 0.89
Min 0 0 -1.96
1st Quartile 0 0 0
Median 0 0 0
3rd Quartile 0 0 0
Max 12.42 12.91 7.11
Notes: This table shows descriptive statistics for the interaction between the Post-1700 indicator variable and the natural logarithm of potato suitable areas for FAO-GAEZ versions 3 and 4 with the low input specification. The unit of observation is the country-year pair. “S.D.” stands for standard deviation.
Table 2 shows the estimated results of Equation (8). Column (1) shows the result as reported by Nunn and Qian (2011), which the authors interpret as an increase of 1% in suitable land for potato increasing population by 0.04% after year 1700. Column (2) is our baseline and replicates Nunn and Qian (2011)’s result using the low input specification of FAO-GAEZ version 3. Reassuringly, the result in Column (2) is very close to the one reported in Column (1), implying that we can work with low input productivity data version 3 in our next steps.
Table 2: The Impact of the Potato in Total Population
log total population
(1) (2) (3) (4) (5) (6)
$I^{Post}\cdot\log(1+A_{i}^{potato})$ 0.043 0.039 0.038 0.043 0.038 [0.029, 0.229]
(0.014) (0.012) (0.013) (0.006) (0.006)
90% CI [0.011, 0.264]
99% CI [0.000, 0.280]
Observations 1,552 1,552 1,552 1,552 1,552 1,552
Clusters 130 130 130 130 130 130
Notes: Observations are at the country-year level. All regressions use a baseline sample of 130 Old World countries. Countries in North and South America are excluded. The periods are 1000, 1100, 1200, 1300, 1400, 1500, 1600, 1700, 1750, 1800, 1850, and 1900. Column (1) reports the result of Nunn and Qian (2011)’s table IV column 5 which uses middle input potato suitability from FAO-GAEZ version 2. To make results more comparable, Columns (2) and (3) report the results for using low input FAO-GAEZ data version 3 and 4, respectively. Columns (4) and (5) show the result when $I^{Post}\cdot\log(1+A_{i}^{potato})$ is normalized to have a variance equal to one. Column (6) presents the result of our partial identification approach. All specifications include year-fixed effects, country-fixed effects, and a range of controls in logs interacted with year dummies: Old World Crops Area, Elevation, Ruggedness, Tropical Area, Maize Area, Silage Maize Area, Sweet Potatoes Area, Cassava Area. Standard errors are clustered at the country level. The confidence interval in Column (6) is built at the 10% and 1% levels.
Column (3) shows the estimate of Equation (8) using FAO-GAEZ version 4. We see this specification as naturally taking place in the coming years, with researchers simply substituting version 3 with version 4 while ignoring that there is information to be extracted by combining both versions. Comparing Columns (2) and (3), we see that changing the version of the FAO-GAEZ data from 3 to 4 results in very mild changes.
We can now discuss the results of our partial identification estimator. The interpretation of these results does not align directly with the results shown so far for two reasons. First, in our approach, we normalize the variance of $I^{Post}\cdot\log(1+A_{i}^{potato})$ such that it equals one (Assumption 2). Second, Nunn and Qian (2011) uses the transformation $\log(1+x)$ which makes it hard to interpret the estimated coefficient as elasticities (Chen and Roth, 2023). For these two reasons, we discuss our results in terms of standard deviations of $I^{Post}\cdot\log(1+A_{i}^{potato})$. To make the comparison easier, Columns (4) and (5) show the result of estimating Equation (8) after normalizing $I^{Post}\cdot\log(1+A_{i}^{potato})$ for each FAO-GAEZ version. In Columns (4) and (5), the interpretation of the results is that an increase of one standard deviation in $I^{Post}\cdot\log(1+A_{i}^{potato})$ causes an increase of 4.3% and 3.8% in population.
Column (6) shows the results associated with our method. As we develop a partial identification strategy, our result is an estimated interval. This interval shows a wide range of possible effects, with an increase of one standard deviation in $I^{Post}\cdot\log(1+A_{i}^{potato})$ leading to an increase in population between 2.9% and 22.9%. Importantly, this interval exhausts all the information contained in the first and second moments of the data-generating process, implying that the large estimate for the upper bound is not a result of our estimator being loose. Such a wide estimated interval, especially in comparison to the small standard errors reported in Columns (4) and (5), suggests that model uncertainty with respect to the existence of non-classical measurement error is more concerning than sample uncertainty.
We also report in Column (6) the confidence intervals at the 10% and 1% levels. Even at the 10%-confidence level, we see that the lower bound of the effect is a quarter of the point-estimate associated with the FAO-GAEZ version 3 (Column (4)). We also highlight that, when ignoring measurement error, Nunn and Qian (2011)’s results and our results in Columns (2)-(5) are all significantly different from zero at the 1%-confidence level. However, when we account for measurement error (Column (6)), we cannot reject that the effect of potatoes introduction on population is zero at the 1%-confidence level.
Bustos et al. (2016)
Bustos et al. (2016) studies the impact of agricultural productivity on structural transformation. The authors focus on the increase of soybean productivity in Brazil after the introduction of genetic modified seeds in 2003. Their empirical specification uses panel data to estimate Equation (9):
$\Delta y_{it}=\alpha_{t}+\beta\cdot\Delta A_{it}^{soy}+\sum_{j=1}^{J}\alpha_{j}\cdot w_{i,j}+\epsilon_{jt},$ (9)
where $y_{it}$ is the employment share of the manufacturing sector in municipality $i$ and year $t$, and $\Delta A_{it}^{soy}$ is the soybean productivity shock, which is the variable of interest $X^{*}$. Bustos et al. (2016) exploits an arguably exogenous shock in soybean productivity: the introduction of genetically modified seeds in 2003, i.e., they use this increase in productivity as an exogenous change in soybean productivity.
The problem is that there is no measure of the increase in productivity generated by genetically modified seeds. Instead, the authors propose to use the difference between soybean productivity for high and low inputs of FAO-GAEZ version 3. This choice naturally raises questions about the existence of measurement error.
We begin by discussing the differences in soybean productivity in versions 3 and 4. We are able to reproduce the numbers of productivity using Bustos et al. (2016)’s replication package. The correlation between the change in productivity variable in the replication package and the one generated by us is of 0.99.151515There are many reasons for differences in the computed productivity. In the replication package, there is no mention of how the variables were transformed from raster and vector formats to the tabular format the authors use. We use the geometry of the minimum comparable areas that is present in the replication package in order to compute the average productivity within each municipality. The replication package does not have the raw soybean productivity data, but we know of no different vintages of the FAO-GAEZ version 3 and, therefore, we have no reason to believe that the data is different. The differences then are likely to occur from different ways of assigning pixels of soybean productivity to the municipality geometry. For example: are all touched pixels considered? If yes, how? What is the projection used to make compatible the raster file of FAO-GAEZ with the vector file of the municipalities? There are also small holes in some municipalities in the Northeast region that we do not know if it was treated somehow after the data were generated. To ensure that all the differences we point out are coming from different FAO-GAEZ versions and not different data handling, we compare the results across our constructed productivity variables.
We focus on the difference between soybean productivity with high and low use of inputs ($\Delta A_{i}^{soy}\coloneqq A^{soy}_{i,high}-A^{soy}_{i,low}$), which is interpreted as the productivity shock brought by the genetically modified seeds. In our setting, the variable $\Delta A_{i}^{soy}$ is the variable of interest $X^{*}$. Table 3 shows descriptive statistics of the changes in productivity. On average, version 4 measures a higher soybean productivity of 170 kg per hectare. Overall, version 4 has lower spatial heterogeneity than version 3.
Table 3: Descriptive Statistics: Change in Soybean Productivity (ton/ha)
$\Delta A_{i}^{soy}$ (ton/ha)
V 3 V 4 V 4 - V 3
(1) (2) (3)
Mean 1.79 1.96 0.17
S.D. 0.83 0.72 0.27
Min -0.26 -0.32 -1.09
1st Quartile 1.17 1.40 0.02
Median 1.82 1.94 0.15
3rd Quartile 2.43 2.49 0.33
Max 4.00 4.00 2.27
Notes: This table shows descriptive statistics for the difference between soybean productivity with high and low input. The first two columns show the statistics for FAO-GAEZ versions 3 and 4. The third column shows statistics for the difference between versions. The productivity is measured in tons per hectare. “S.D.” stands for standard deviation.
Table 4 shows the estimated results of Equation (9). We focus on Column (2) of Table 9 of the original article, which reports the impact of agricultural productivity on employment share in manufacturing. Column (1) in Table 4 shows the result as reported by Bustos et al. (2016), which shows that an increase of one standard deviation in potential soy productivity leads to an increase of 1.7 p.p $0.83\times 0.021$] on the share of manufacturing employment.
Table 4: The Effect of Agricultural Technological Change on Manufacturing
$\Delta$ Employment share
(1) (2) (3) (4) (5) (6)
$\Delta$ soy productivity 0.021 0.022 0.029 0.010 0.014 [0.012, 0.055]
(0.002) (0.002) (0.002) (0.001) (0.001)
90% CI [0.010, 0.058]
99% CI [0.009, 0.059]
Observations 4,149 4,149 4,149 4,149 4,149 4,149
Clusters 4,149 4,149 4,149 4,149 4,149 4,149
Notes: Observations are at the municipality level. All results use a baseline sample of 4,149 municipalities. Changes in dependent variables are calculated over the years 2000 and 2010. All municipality controls are from the population census of 1991: share rural population, log of income per capita, log of population density, literacy rate. Column (1) reports the result of Bustos et al. (2016)’s Column 2 in Table 9, which uses soybean suitability from FAO-GAEZ version 3. To make results more comparable, Columns (2) and (3) report our results using FAO-GAEZ data versions 3 and 4, respectively. Columns (4) and (5) show the result when $\Delta A_{it}^{soy}$ is normalized to have variance equal to one. Column (6) presents our partial identification approach. Standard errors are robust to heteroskedasticity. Confidence intervals in Column (6) are built at the 10% and 1% levels.
In Columns (2) and (3), we replicate Bustos et al. (2016)’s result using the variables we constructed based on FAO-GAEZ versions 3 and 4. Reassuringly, the result in Column (2) is very close to the one reported in Column (1), since they both use version 3 of the explanatory variable. In Column (3), we estimate Equation (9) using FAO-GAEZ version 4. Comparing Columns (2) and (3), we see a significant increase in the estimated effect when we substitute version 3 of the FAO-GAEZ data for version 4.
We can now discuss the results of our proposed econometric tool. To make the comparison easier, Columns (4) and (5) show the result of estimating Expression (9) with FAO-GAEZ versions 3 and 4 after normalizing the main explanatory variable. In Columns (4) and (5), the interpretation of the estimated results is that an increase of one standard deviation in $\Delta A_{it}^{soy}$ causes an increase of 1 p.p and 1.4 p.p in the share of manufacturing jobs.
Column (6) shows the results associated with our method. The estimated interval shows a wide range of possible effects, but the lower bound is very close to the point estimates associated with OLS regressions using either proxy variable (Columns (4) and (5)). This finding suggests that measurement error may not be a critical issue in this empirical application.
We also report in Column (6) the confidence intervals at the 10% and 1% levels. Even at the 10%-confidence level, we see that the lower bound is very close to the point estimates associated with OLS regressions using FAO-GAEZ versions 3 and 4 (Columns (4) and (5)). Additionally, when ignoring measurement error, Bustos et al. (2016)’s result and our results in Columns (2)-(5) are all significant at the 1% level. Interestingly, when we account for measurement error, our partial identification approach still rejects the null of no effect at the 1%-confidence level. This results suggests that Bustos et al. (2016)’s result is robust to measurement error.
Acharya et al. (2016)
Acharya et al. (2016) study the impact of slavery on contemporary political attitudes in the South of the United States. In particular, the authors estimate the following equation:
$y_{i}=\beta\cdot S_{i}+\sum_{j=1}^{J}\alpha_{j}\cdot w_{i,j}+\epsilon_{j},$
where $y_{i}$ is the proportion of republicans in the white population in county $i$ pooled from Cooperative Congressional Election Study for the years 2006, 2008, 2009, 2010 and 2011, and $S_{i}$ is the proportion of slaves in county $i$ in 1860. The authors point out at least two potential problems with estimating the above expression via OLS: measurement error in the explanatory variable and omitted variable bias. To overcome these concerns the authors propose to use the suitability of cotton from FAO-GAEZ data version 3 as an instrument for slavery $(S_{i})$. To fit this approach within our framework, we focus on the reduced-form specification given by Equation (10):
$y_{i}=\beta\cdot A_{i}^{cotton}+\sum_{j=1}^{J}\alpha_{j}\cdot W_{i,j}+\epsilon_{i}$ (10)
The authors use an average of the FAO-GAEZ version 3 with high and intermediate inputs. As FAO-GAEZ version 4 has no intermediate input specification, we reproduce their paper using the high input specification. In our setting, the variable $A_{i}^{cotton}$ is the variable of interest $X^{*}$.
We start by discussing the differences in FAO-GAEZ cotton suitability in versions 3 and 4. Table 5 shows that, for the counties in the South of the United States, cotton productivity is lower and less heterogeneous in version 4 than in version 3.
Table 5: Descriptive Statistics: Cotton Productivity (ton/ha)
$A_{i}^{cotton}$ (ton/ha)
V 3 V 4 V 4 - V 3
(1) (2) (3)
Mean 0.56 0.45 -0.11
S.D. 0.27 0.21 0.12
Min 0.00 0.00 -0.66
1st Quartile 0.42 0.31 -0.17
Median 0.62 0.50 -0.12
3rd Quartile 0.77 0.62 -0.04
Max 1.17 0.95 0.52
Notes: This table shows descriptive statistics for the cotton productivity with high input. The first two columns show the statistics for FAO-GAEZ versions 3 and 4. The third column shows statistics for the difference between versions. The productivity is measured in tons per hectare. “S.D.” stands for standard deviation.
In Table 6, we estimate Equation (10) with the FAO-GAEZ data versions 3 and 4. Once more, to make results comparable with our partial identification approach, we normalize the $A^{cotton}_{j}$ variable to have a variance equal to one. The interpretation is, then, that an increase of one standard deviation in cotton suitability generates an increase of the share of Republicans in the White population of 2 p.p using FAO-GAEZ version 3 (Column (1)) and a non-significant effect of $0.9$ p.p using FAO-GAEZ version 4 (Column (2)). Such a large difference between these point estimates suggests the importance of accounting for measurement error in this application.
Table 6: The Effect of Cotton Productivity on the Proportion of Republicans in the White Population
Proportion of Republicans
(1) (2) (3)
Cotton Productivity 0.020 0.009 [0.007, 0.245]
(0.008) (0.007)
90% CI [0.000, 0.257]
99% CI [0.000, 0.261]
Observations 1,180 1,180 1,180
Clusters 1,180 1,180 1,180
Notes: Observations are at the county level. All results include state fixed effects and the following controls as measured in 1860: log of the total county population, proportion of farms in the county smaller than 50 acres, inequality of farmland holdings as measured by the Gini coefficient for landownership, log of total farm value per improved acre of farmland in the county, log of the acres of improved farmland; the proportion of total population that is free black, access to rails and steamboat-navigable rivers or canals, log of the county acreage, ruggedness of the county, latitude and longitude as well as their squared terms. We keep the weighting scheme of the original paper with the within-county sample size (appropriately weighted by the sampling weights) as weights. Columns (1) and (2) report the results of the reduced form counterpart of Acharya et al. (2016)’s Column 2 in Table 2 using cotton suitability from FAO-GAEZ versions 3 and 4, respectively. Column (3) presents results associated with our partial identification approach. Standard errors are robust to heteroskedasticity. Confidence intervals in Column (3) are built at the 10% and 1% levels.
Column (3) shows the results associated with our method. The estimated interval accounts for nonclassical measurement error in the explanatory variable and shows a wide range of possible effects. Interestingly, the estimated lower bound is almost a third of the point estimate associated with FAO-GAEZ version 3 (Column (1)) and is even lower than the point estimate associated with FAO-GAEZ version 4 (Column (4)), suggesting that accounting for measurement error in this empirical context is relevant. Lastly, when we account for sample uncertainty when estimating our bounds in Column (3), we do not reject the null of no effect at the 10%-confidence level. [^16] The estimated 90%-confidence interval is $\left[-0.006,0.257\right]$ while the 99%-confidence interval is $\left[-0.013,0.261\right]$. In Column (3) of Table 6, we report the intersection of these estimated intervals with $\mathbb{R}_{+}$ due to Assumption 3.
Recommendations to the Practitioner
In this section, we provide a detailed, 5-step guide on how to use our method in any empirical application with a mismeasured treatment variable and two proxies.
First, the practitioner must reflect on the exogeneity of the correctly measured treatment variable (Assumption 1). If the true treatment were observable, could the practitioner identify the treatment effect of interest using a linear regression model? This type of argument is present in any empirical application and practitioners still need to think about it when using our proposed method. If the answer is positive, proceed to the next step. If the answer is negative, the practitioner should use a different approach (e.g., instrumental variables).
Second, the practitioner must reflect on the exogeneity of the measurement errors (Assumption 4). Are the measurement error terms uncorrelated with unobserved factors in the outcome equation? To answer this question, it is important to think about the causes of the measurement error. Is it due to arbitrary modeling decisions about imputation or interpolation methods? Or is it due to better institutions being able to collect better data? If the first scenario is more likely, then Assumption 4 is plausible. If the practitioner believes that Assumption 4 is valid, proceed to the next step.
Third, the practitioner must reflect on their priors about the sign of the coefficient of interest (Assumption 3) and about the sign of the correlation between the measurement error terms (Assumption 5).
To think about the sign of the coefficient of interest, the practitioner may compare different economic theories or analyze previous empirical results using different datasets. Do they agree about the direction of the true effect? If so, Assumption 3 is plausible.
To think about the sign of the correlation between the measurement error terms, the practitioner must, once more, think about the causes of the measurement error. If it is hard to collect high-quality data on some regions, then Assumption 5 is plausible. If measurement error is due to different modeling choices, then Assumption 5 may not be plausible.
  1. 1.
    If the practitioner believes that Assumptions 3 and 5 are plausible, then proceed with the method described in Section 3.
  2. 2.
    If the practitioner believes that only Assumption 3 is plausible, then proceed with the method described in Appendix C.
  3. 3.
    If the practitioner believes that only Assumption 5 is plausible, then proceed with the method described in Appendix D.
Fourth, the practitioner must estimate the bounds around the true coefficient of interest using the estimator proposed in Appendix E.1. The estimated bounds are interesting for at least three aspects:
  1. 1.
    If all three terms in the estimated lower bound are similar to each other, then the results are robust to measurement error, and the practitioner may focus their analysis on their OLS results. This scenario is illustrated by our analysis of the dataset used by Bustos et al. (2016).
  2. 2.
    If the estimated lower bound is numerically close to zero, then the true coefficient of interest may be small. This type of result suggests that the economic relevance of the target parameter may not be large. This scenario is illustrated by our analysis of the dataset used by Acharya et al. (2016).
  3. 3.
    If the estimated interval is wide, then the model uncertainty generated by the possible existence of measurement error is relevant. This scenario is illustrated by our analysis of the dataset used by Nunn and Qian (2011).
Lastly, the practitioner must estimate the confidence intervals around the true coefficient of interest using the algorithm proposed in Appendix E.2. If the confidence interval intersects zero, then the “null hypothesis that the true coefficient of interest equals zero” is not rejected. In this case, the data does not provide evidence that there is an effect of $X^{*}$ on $Y$ when we account for non-classical measurement error.
Conclusion
In this paper, we develop a novel partial identification analytical method to identify the effect of a mismeasured variable when two proxies are available. Importantly, our estimator relies on assumptions no stronger than those commonly applied in linear models that ignore measurement error. Moreover, our partial identification approach provides bounds that (i) incorporate all available information from the first two moments of the data distribution and (ii) compare favorably against alternative approaches, such as a simple reduced-form estimand and the method proposed by Lubotsky and Wittenberg (2006).
We further explore the implications of relaxing our assumptions. In particular, we relax restrictions on the sign of the target parameter (Appendix D) and the correlation of measurement errors (Appendix C), deriving wider identified intervals in both cases. Thus, modeling assumptions should be chosen in a way that best fits each application.
To illustrate our methodology, we reevaluate the results of three empirical studies: Nunn and Qian (2011), Bustos et al. (2016) and Acharya et al. (2016). To do so, we replicate their findings using the reduced-form approach, which ignores measurement error, and compare these results with those obtained through our partial identification method. Although we find that the results by Bustos et al. (2016) are robust to nonclassical measurement error, our method reaches very different conclusions when compared to the findings of Nunn and Qian (2011) and Acharya et al. (2016). In particular, we find that the impact of agricultural productivity may be smaller than previously reported.
Appendix A Proofs of the Main Results
Proof of Proposition 1
We break the proof of Proposition 1 in 4 parts:
  1. 1.
    Section A.1.1 derives the restrictions imposed by the first two moments of the observable data on the underlying data-generating process.
  2. 2.
    Section A.1.2 shows that $LB\leq\beta\leq UB$ if $UB\geq LB$.
  3. 3.
    Section A.1.3 shows that, if $LB>UB$, then there is no data-generating process that satisfies Assumptions 1-6 and the restrictions imposed by the first two moments of the data distribution.
  4. 4.
    Section A.1.4 defines the meaning of exhausting all the information contained in the first two moments of the data distribution and shows that, if $UB\geq LB$, then the interval $\left[LB,UB\right]$ exhausts all the information contained in the first two moments of the data distribution.
A.1.1 Data Restrictions
In this section, we derive the restrictions imposed by the first two moments of the observable data distribution $\left(Y_{res},Z_{1,res},Z_{2,res}\right)$ on the distribution of the latent variables $\left(X_{res}^{*},\epsilon_{res},U_{1},U_{2}\right)$. First, we state all nine data restrictions in a lemma. Afterwards, we prove them.
Lemma A.1 (Data Restrictions)
Under Assumptions 1, 2 and 4, we have that
$\mathbb{E}\left[Y_{res}\right]$ $=0$ (A.1)
$\mathbb{E}\left[Z_{1,res}\right]$ $=\mathbb{E}\left[U_{1}\right]$ (A.2)
$\mathbb{E}\left[Z_{2,res}\right]$ $=\mathbb{E}\left[U_{2}\right]$ (A.3)
$Var\left(Y_{res}\right)$ $=\beta^{2}+Var\left(\epsilon_{res}\right)$ (A.4)
$Var\left(Z_{1,res}\right)$ $=1+Var\left(U_{1}\right)+2\cdot Cov\left(X_{res}^{*},U_{1}\right)$ (A.5)
$Var\left(Z_{2,res}\right)$ $=1+Var\left(U_{2}\right)+2\cdot Cov\left(X_{res}^{*},U_{2}\right)$ (A.6)
$Cov\left(Z_{1,res},Y_{res}\right)$ $=\beta+\beta\cdot Cov\left(X_{res}^{*},U_{1}\right)$ (A.7)
$Cov\left(Z_{2,res},Y_{res}\right)$ $=\beta+\beta\cdot Cov\left(X_{res}^{*},U_{2}\right)$ (A.8)
$Cov\left(Z_{1,res},Z_{2,res}\right)$ $=1+Cov\left(X_{res}^{*},U_{1}\right)+Cov\left(X_{res}^{*},U_{2}\right)+Cov\left(U_{1},U_{2}\right).$ (A.9)
Proof.
To prove Equation (A.1), note that:
$\mathbb{E}\left[Y_{res}\right]$ $=\mathbb{E}\left[\beta\cdot X_{res}^{*}+\epsilon_{res}\right]$ by Equation (2)
$=\beta\cdot\mathbb{E}\left[X_{res}^{*}\right]+\mathbb{E}\left[\epsilon_{res}\right]$ by linearity of the expectation operator
$=0$ $\text{because }X_{res}^{*}\text{ and }\epsilon_{res}\text{ are residuals.}$
To prove Equations (A.2) and (A.3), fix $k\in\left\{1,2\right\}$ arbitrarily and observe that
$\mathbb{E}\left[Z_{k,res}\right]$ $=\mathbb{E}\left[X_{res}^{*}+U_{k}\right]$ by Equation (3)
$=\mathbb{E}\left[X_{res}^{*}\right]+\mathbb{E}\left[U_{k}\right]$ by linearity of the expectation operator
$=\mathbb{E}\left[U_{k}\right]$ $\text{because }X_{res}^{*}\text{ is a residual.}$
To prove Equation (A.4), notice that:
$Var\left(Y_{res}\right)$ $=Var\left(\beta\cdot X_{res}^{*}+\epsilon_{res}\right)$
by Equation (2)
$=\beta^{2}\cdot Var\left(X_{res}^{*}\right)+Var\left(\epsilon_{res}\right)+2\cdot\beta\cdot Cov\left(X_{res}^{*},\epsilon_{res}\right)$
by the properties of the variance
$=\beta^{2}+Var\left(\epsilon_{res}\right)$
by Assumptions 1 and 2.
To prove Equations (A.5) and (A.6), fix $k\in\left\{1,2\right\}$ arbitrarily and observe that
$Var\left(Z_{k,res}\right)$ $=Var\left(X_{res}^{*}+U_{k}\right)$
by Equation (3)
$=Var\left(X_{res}^{*}\right)+Var\left(U_{k}\right)+2\cdot Cov\left(X_{res}^{*},U_{k}\right)$
by the properties of the variance
$=1+Var\left(U_{k}\right)+2\cdot Cov\left(X_{res}^{*},U_{k}\right)$
by Assumption 2.
To prove Equations (A.7) and (A.8), fix $k\in\left\{1,2\right\}$ arbitrarily and note that
$Cov\left(Z_{k,res},Y_{res}\right)$
$\hskip 20.0pt=Cov\left(X_{res}^{*}+U_{k},\beta\cdot X_{res}^{*}+\epsilon_{res}\right)$
by Equations (2) and (3)
$\hskip 20.0pt=\beta\cdot Var\left(X_{res}^{*}\right)+Cov\left(X_{res}^{*},\epsilon_{res}\right)+\beta\cdot Cov\left(X_{res}^{*},U_{k}\right)+Cov\left(U_{k},\epsilon_{res}\right)$
by the properties of the covariance
$\hskip 20.0pt=\beta+\beta\cdot Cov\left(X_{res}^{*},U_{k}\right)$
by Assumptions 1, 2 and 4.
To prove Equation (A.9), observe that:
$Cov\left(Z_{1,res},Z_{2,res}\right)$
$\hskip 20.0pt=Cov\left(X_{res}^{*}+U_{1},X_{res}^{*}+U_{2}\right)$
by Equation (3)
$\hskip 20.0pt=Var\left(X_{res}^{*}\right)+Cov\left(X_{res}^{*},U_{1}\right)+Cov\left(X_{res}^{*},U_{2}\right)+Cov\left(U_{1},U_{2}\right)$
by the properties of the covariance
$\hskip 20.0pt=1+Cov\left(X_{res}^{*},U_{1}\right)+Cov\left(X_{res}^{*},U_{2}\right)+Cov\left(U_{1},U_{2}\right)$
by Assumption 2.
A.1.2 $LB\leq\beta\leq UB$ if $UB\geq LB$
In this section, we consider the case when $UB\geq LB$ and prove that (i) $\beta\leq UB$, (ii) $\beta\geq LB$ when $\beta=0$, and (iii) $\beta\geq LB$ when $\beta>0$. In this proof, we do not impose that $LB\geq 0$.
Part 1: $\beta\leq UB$
Rearranging Equation (A.4), we have that
$\beta^{2}$ $=Var\left(Y_{res}\right)-Var\left(\epsilon_{res}\right)$
$\leq Var\left(Y_{res}\right)$
because $Var\left(\epsilon_{res}\right)\geq 0$, implying that
$\beta\leq\sqrt{Var\left(Y_{res}\right)}=UB$
by Assumption 3.
Part 2: $\beta\geq LB$ when $\beta=0$
By combining Equations (A.7) and (A.8) with $\beta=0$, we find that
$Cov\left(Z_{1,res},Y_{res}\right)=0$
and
$Cov\left(Z_{2,res},Y_{res}\right)=0,$
implying that
$LB=0,$
according to Equation (4), and that
$\beta=0=LB.$
Part 3: $\beta\geq LB$ when $\beta>0$
Having in mind that $\beta>0$ and rearranging Equations (A.7) and (A.8), we have that
$Cov\left(X_{res}^{*},U_{1}\right)=\dfrac{Cov\left(Z_{1,res},Y_{res}\right)}{\beta}-1$ (A.10)
and
$Cov\left(X_{res}^{*},U_{2}\right)=\dfrac{Cov\left(Z_{2,res},Y_{res}\right)}{\beta}-1.$ (A.11)
Rearranging Equation (A.5), we find that
$Var\left(U_{1}\right)$ $=Var\left(Z_{1,res}\right)-1-2\cdot Cov\left(X_{res}^{*},U_{1}\right)$
$=Var\left(Z_{1,res}\right)-1-2\cdot\left[\dfrac{Cov\left(Z_{1,res},Y_{res}\right)}{\beta}-1\right]$
according to Equation (A.10)
$=1+Var\left(Z_{1,res}\right)-2\cdot\dfrac{Cov\left(Z_{1,res},Y_{res}\right)}{\beta}$
Since $Var\left(U_{1}\right)\geq 0$ and $\beta>0$, the last equation implies that
$\begin{array}[]{crcl}&1+Var\left(Z_{1,res}\right)-2\cdot\dfrac{Cov\left(Z_{1,res},Y_{res}\right)}{\beta}&\geq&0\\ \Leftrightarrow&1+Var\left(Z_{1,res}\right)&\geq&2\cdot\dfrac{Cov\left(Z_{1,res},Y_{res}\right)}{\beta}\\ \Leftrightarrow&\beta\cdot\left(1+Var\left(Z_{1,res}\right)\right)&\geq&2\cdot Cov\left(Z_{1,res},Y_{res}\right),\end{array}$
implying that
$\beta\geq 2\cdot\dfrac{Cov\left(Z_{1,res},Y_{res}\right)}{1+Var\left(Z_{1,res}\right)}.$ (A.12)
Combining Equations (A.6) and (A.11) and using the same steps used to derive Equation (A.12), we have that
$\beta\geq 2\cdot\dfrac{Cov\left(Z_{2,res},Y_{res}\right)}{1+Var\left(Z_{2,res}\right)}.$ (A.13)
Substituting Equations (A.10) and (A.11) in Equation (A.9), we have that
$Cov\left(Z_{1,res},Z_{2,res}\right)$ $=1+Cov\left(X_{res}^{*},U_{1}\right)+Cov\left(X_{res}^{*},U_{2}\right)+Cov\left(U_{1},U_{2}\right)$
$=1+\left[\dfrac{Cov\left(Z_{1,res},Y_{res}\right)}{\beta}-1\right]+\left[\dfrac{Cov\left(Z_{2,res},Y_{res}\right)}{\beta}-1\right]+Cov\left(U_{1},U_{2}\right)$
$=-1+\dfrac{Cov\left(Z_{1,res},Y_{res}\right)+Cov\left(Z_{2,res},Y_{res}\right)}{\beta}+Cov\left(U_{1},U_{2}\right)$
$\geq-1+\dfrac{Cov\left(Z_{1,res},Y_{res}\right)+Cov\left(Z_{2,res},Y_{res}\right)}{\beta}$
by Assumption 5
implying that
$\beta\geq\dfrac{Cov\left(Z_{1,res},Y_{res}\right)+Cov\left(Z_{2,res},Y_{res}\right)}{1+Cov\left(Z_{1,res},Z_{2,res}\right)}$ (A.14)
because $\beta>0$ and $Cov\left(Z_{1,res},Z_{2,res}\right)>-1$ according to Assumption 6.
Combining Equations (4) and (A.12)-(A.14), we have that
$\beta\geq LB.$
A.1.3 No Valid Data-Generating Process if $LB>UB$
In this section, we show that, when $LB>UB$, there is no combination of $\beta\geq 0$ and distribution of latent variables $\left(X_{res}^{*},\epsilon_{res},U_{1},U_{2}\right)$ that satisfies Assumptions 1-6 and the data restrictions described in Lemma A.1. We prove this result by contradiction.
Assume that $LB>UB$ and that there exist parameter $\beta\geq 0$ and variables $\left(X_{res}^{*},\epsilon_{res},U_{1},U_{2}\right)$ such that Assumptions 1-6 and Equations (A.1)-(A.9) are valid.
Following the same steps described in Parts 2 and 3 in Section A.1.2, we must have that $\beta\geq LB$. Rearranging Equation (A.4), we have that
$Var\left(\epsilon_{res}\right)$ $=Var\left(Y_{res}\right)-\beta^{2}$
$\leq Var\left(Y_{res}\right)-\left(LB\right)^{2}$
[…]
$=\left(UB\right)^{2}-\left(LB\right)^{2}$
according to Equation (5)
$<0\hskip 10.0pt\text{because }LB>UB,$
which is a contradiction.
A.1.4 $\left[LB,UB\right]$ exhausts all the information contained in the first two moments of the data distribution if $UB\geq LB$
First, we define the meaning of exhausting all the information contained in the first two moments of the data distribution. Second, we show that, if $UB\geq LB$, then the interval $\left[LB,UB\right]$ exhausts all the information contained in the first two moments of the data distribution.
Definition A.1
A set $\mathcal{B}\subseteq\mathbb{R}$ exhausts all the information contained in the first two moments of the data distribution if, for any candidate target parameter $\tilde{\beta}\in\mathcal{B}$, there exist candidate latent variables $\left(\tilde{X}_{res}^{*},\tilde{\epsilon}_{res},\tilde{U}_{1},\tilde{U}_{2}\right)$ and candidate observable variables $\left(\tilde{Y}_{res},\tilde{Z}_{1,res},\tilde{Z}_{1,res}\right)$ such that
  1. 1.
    the candidate observable variables are induced by the candidate latent variables according to the model restrictions imposed by Equations (2) and (3), i.e.,
    $\tilde{Y}_{res}$ $\coloneqq\tilde{\beta}\cdot\tilde{X}_{res}^{*}+\tilde{\epsilon}_{res}$ (A.15)
    $\tilde{Z}_{1,res}$ $\coloneqq\tilde{X}_{res}^{*}+\tilde{U}_{1}$ (A.16)
    $\tilde{Z}_{2,res}$ $\coloneqq\tilde{X}_{res}^{*}+\tilde{U}_{2};$ (A.17)
  2. 2.
    the candidate latent variables and the candidate target parameter satisfy the model restrictions imposed by Assumptions 1-5, i.e.,
    $Cov\left(\tilde{X}_{res}^{*},\tilde{\epsilon}_{res}\right)$ $=0$ (A.18)
    $Var\left(\tilde{X}_{res}^{*}\right)$ $=1$ (A.19)
    $\tilde{\beta}$ $\geq 0$ (A.20)
    $Cov\left(\tilde{U}_{1},\tilde{\epsilon}_{res}\right)$ $=0$ (A.21)
    $Cov\left(\tilde{U}_{2},\tilde{\epsilon}_{res}\right)$ $=0$ (A.22)
    $Cov\left(\tilde{U}_{1},\tilde{U}_{2}\right)$ $\geq 0$ (A.23)
  3. 3.
    the first two moments of the candidate observable variables are equal to the first two moments of the true observable variables, i.e.,
    $\mathbb{E}\left[\tilde{Y}_{res}\right]$ $=\mathbb{E}\left[Y_{res}\right]$ (A.24)
    $\mathbb{E}\left[\tilde{Z}_{1,res}\right]$ $=\mathbb{E}\left[Z_{1,res}\right]$ (A.25)
    $\mathbb{E}\left[\tilde{Z}_{2,res}\right]$ $=\mathbb{E}\left[Z_{2,res}\right]$ (A.26)
    $Var\left(\tilde{Y}_{res}\right)$ $=Var\left(Y_{res}\right)$ (A.27)
    $Var\left(\tilde{Z}_{1,res}\right)$ $=Var\left(Z_{1,res}\right)$ (A.28)
    $Var\left({Z}_{2,res}\right)$ $=Var\left(Z_{2,res}\right)$ (A.29)
    $Cov\left(\tilde{Z}_{1,res},\tilde{Y}_{res}\right)$ $=Cov\left(Z_{1,res},Y_{res}\right)$ (A.30)
    $Cov\left(\tilde{Z}_{2,res},\tilde{Y}_{res}\right)$ $=Cov\left(Z_{2,res},Y_{res}\right)$ (A.31)
    $Cov\left(\tilde{Z}_{1,res},\tilde{Z}_{2,res}\right)$ $=Cov\left(Z_{1,res},Z_{2,res}\right).$ (A.32)
Second, we show that, if $UB\geq LB\geq 0$, then the interval $\left[LB,UB\right]$ exhausts all the information contained in the first two moments of the data distribution.
Assume that $UB\geq LB$ and fix $\tilde{\beta}\in\left[LB,UB\right]$ arbitrarily. We consider two cases: (i) $\tilde{\beta}>0$, and (ii) $\tilde{\beta}=0$. In both cases, the proof is by construction.
Part 1: $\tilde{\beta}>0$
We need to create a distribution for the candidate latent variables $\left(\tilde{X}_{res}^{*},\tilde{\epsilon}_{res},\tilde{U}_{1},\tilde{U}_{2}\right)$. The candidate observable variables $\left(\tilde{Y}_{res},\tilde{Z}_{1,res},\tilde{Z}_{1,res}\right)$ are created based on Equations (A.15)-(A.17).
We impose
$\left(\tilde{X}_{res}^{*},\tilde{\epsilon}_{res},\tilde{U}_{1},\tilde{U}_{2}\right)\sim N\left(\boldsymbol{\mu},\boldsymbol{\Omega}\right),$
i.e., $\left(\tilde{X}_{res}^{*},\tilde{\epsilon}_{res},\tilde{U}_{1},\tilde{U}_{2}\right)$ are jointly normally distributed with mean $\boldsymbol{\mu}$ and covariance matrix $\boldsymbol{\Omega}$.
The vector $\boldsymbol{\mu}$ is given by
$\mathbb{E}\left[\tilde{X}_{res}^{*}\right]$ $\coloneqq 0$ (A.33)
$\mathbb{E}\left[\tilde{\epsilon}_{res}\right]$ $\coloneqq 0$ (A.34)
$\mathbb{E}\left[\tilde{U}_{1}\right]$ $\coloneqq\mathbb{E}\left[Z_{1,res}\right]$ (A.35)
$\mathbb{E}\left[\tilde{U}_{1}\right]$ $\coloneqq\mathbb{E}\left[Z_{2,res}\right]$ (A.36)
and the covariance matrix $\boldsymbol{\Omega}$ is given by
$Var\left(\tilde{X}_{res}^{*}\right)$ $\coloneqq 1$ (A.37)
$Var\left(\tilde{\epsilon}_{res}\right)$ $\coloneqq Var\left(Y_{res}\right)-\tilde{\beta}^{2}$ (A.38)
$Var\left(\tilde{U}_{1}\right)$ $\coloneqq 1+Var\left(Z_{1,res}\right)-2\cdot\dfrac{Cov\left(Z_{1,res},Y_{res}\right)}{\tilde{\beta}}$ (A.39)
$Var\left(\tilde{U}_{2}\right)$ $\coloneqq 1+Var\left(Z_{2,res}\right)-2\cdot\dfrac{Cov\left(Z_{2,res},Y_{res}\right)}{\tilde{\beta}}$ (A.40)
$Cov\left(\tilde{X}_{res}^{*},\tilde{\epsilon}_{res}\right)$ $\coloneqq 0$ (A.41)
$Cov\left(\tilde{U}_{1},\tilde{\epsilon}_{res}\right)$ $\coloneqq 0$ (A.42)
$Cov\left(\tilde{U}_{2},\tilde{\epsilon}_{res}\right)$ $\coloneqq 0$ (A.43)
$Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{1}\right)$ $\coloneqq\dfrac{Cov\left(Z_{1,res},Y_{res}\right)}{\tilde{\beta}}-1$ (A.44)
$Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{2}\right)$ $\coloneqq\dfrac{Cov\left(Z_{2,res},Y_{res}\right)}{\tilde{\beta}}-1$ (A.45)
$Cov\left(\tilde{U}_{1},\tilde{U}_{2}\right)$ $\coloneqq 1+Cov\left(Z_{1,res},Z_{2,res}\right)-\left[\dfrac{Cov\left(Z_{1,res},Y_{res}\right)+Cov\left(Z_{2,res},Y_{res}\right)}{\tilde{\beta}}\right$] (A.46)
First, we must show that the above distribution is a valid normal distribution, i.e., we must show that $Var\left(\tilde{\epsilon}_{res}\right)\geq 0$, $Var\left(\tilde{U}_{1}\right)\geq 0$ and $Var\left(\tilde{U}_{2}\right)\geq 0$.
Note that
$Var\left(\tilde{\epsilon}_{res}\right)$ $\coloneqq Var\left(Y_{res}\right)-\tilde{\beta}^{2}$
according to Equation (A.38)
$\geq Var\left(Y_{res}\right)-\left(UB\right)^{2}$
$\hskip 20.0pt\text{because }\tilde{\beta}\leq UB$
$=Var\left(Y_{res}\right)-\left(\sqrt{Var\left(Y_{res}\right)}\right)^{2}$
according to Equation (5)
$=0.$
Now, we show that $Var\left(\tilde{U}_{1}\right)\geq 0$. To do so, we analyze two cases. If $Cov\left(Z_{1,res},Y_{res}\right)=0$, then $Var\left(\tilde{U}_{1}\right)=1+Var\left(Z_{1,res}\right)>0$ according to Equation (A.39). If $Cov\left(Z_{1,res},Y_{res}\right)>0$, then
$Var\left(\tilde{U}_{1}\right)$ $=1+Var\left(Z_{1,res}\right)-2\cdot\dfrac{Cov\left(Z_{1,res},Y_{res}\right)}{\tilde{\beta}}$
according to Equation (A.39)
$\geq 1+Var\left(Z_{1,res}\right)-2\cdot\dfrac{Cov\left(Z_{1,res},Y_{res}\right)}{\nicefrac{{2\cdot Cov\left(Z_{1,res},Y_{res}\right)}}{{\left(1+Var\left(Z_{1,res}\right)\right)}}}$
[…]
$=0.$
We can show that $Var\left(\tilde{U}_{2}\right)\geq 0$ using Equation (A.40) and the steps above.
Hence, we can conclude that the proposed distribution for $\left(\tilde{X}_{res}^{*},\tilde{\epsilon}_{res},\tilde{U}_{1},\tilde{U}_{2}\right)$ is a valid normal distribution.
Second, we must show that the candidate latent variables satisfy Equations (A.18)-(A.23). Since Restrictions (A.18)-(A.22) are satisfied by construction, we only have to show that Restriction (A.23) is satisfied. To do so, we analyze two cases. If $Cov\left(Z_{1,res},Y_{res}\right)+Cov\left(Z_{2,res},Y_{res}\right)=0$, then $Cov\left(\tilde{U}_{1},\tilde{U}_{2}\right)=1+Cov\left(Z_{1,res},Z_{2,res}\right)>0$ according to Equation (A.46) and Assumption 6. If $Cov\left(Z_{1,res},Y_{res}\right)+Cov\left(Z_{2,res},Y_{res}\right)>0$, then
$Cov\left(\tilde{U}_{1},\tilde{U}_{2}\right)$
$\hskip 20.0pt=1+Cov\left(Z_{1,res},Z_{2,res}\right)-\left[\dfrac{Cov\left(Z_{1,res},Y_{res}\right)+Cov\left(Z_{2,res},Y_{res}\right)}{\tilde{\beta}}\right$]
according to Equation (A.46)
$\hskip 20.0pt\geq 1+Cov\left(Z_{1,res},Z_{2,res}\right)-\left[\dfrac{Cov\left(Z_{1,res},Y_{res}\right)+Cov\left(Z_{2,res},Y_{res}\right)}{\nicefrac{{\left(Cov\left(Z_{1,res},Y_{res}\right)+Cov\left(Z_{2,res},Y_{res}\right)\right)}}{{\left(1+Cov\left(Z_{1,res},Z_{2,res}\right)\right)}}}\right]$
[…]
$\hskip 20.0pt=0,$
implying that all the model restrictions are satisfied.
Lastly, we must show that the data restrictions (Equations (A.24)-(A.32)) are valid. Since we have shown that candidate latent variables satisfy Equations (A.18)-(A.23), we can follow the same steps used to prove Lemma A.1 to show that
$\mathbb{E}\left[\tilde{Y}_{res}\right]$ $=0$ (A.47)
$\mathbb{E}\left[\tilde{Z}_{1,res}\right]$ $=\mathbb{E}\left[\tilde{U}_{1}\right]$ (A.48)
$\mathbb{E}\left[\tilde{Z}_{2,res}\right]$ $=\mathbb{E}\left[\tilde{U}_{2}\right]$ (A.49)
$Var\left(\tilde{Y}_{res}\right)$ $=\tilde{\beta}^{2}+Var\left(\tilde{\epsilon}_{res}\right)$ (A.50)
$Var\left(\tilde{Z}_{1,res}\right)$ $=1+Var\left(\tilde{U}_{1}\right)+2\cdot Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{1}\right)$ (A.51)
$Var\left(\tilde{Z}_{2,res}\right)$ $=1+Var\left(\tilde{U}_{2}\right)+2\cdot Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{2}\right)$ (A.52)
$Cov\left(\tilde{Z}_{1,res},\tilde{Y}_{res}\right)$ $=\tilde{\beta}+\tilde{\beta}\cdot Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{1}\right)$ (A.53)
$Cov\left(\tilde{Z}_{2,res},\tilde{Y}_{res}\right)$ $=\tilde{\beta}+\tilde{\beta}\cdot Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{2}\right)$ (A.54)
$Cov\left(\tilde{Z}_{1,res},\tilde{Z}_{2,res}\right)$ $=1+Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{1}\right)+Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{2}\right)+Cov\left(\tilde{U}_{1},\tilde{U}_{2}\right).$ (A.55)
Based on Equation (A.47), we have that
$\mathbb{E}\left[\tilde{Y}_{res}\right]$ $=0=\mathbb{E}\left[Y_{res}\right]$
according to Equation (A.1), implying that Equation (A.24) is valid.
Based on Equation (A.48), we have that
$\mathbb{E}\left[\tilde{Z}_{1,res}\right]$ $=\mathbb{E}\left[\tilde{U}_{1}\right]=\mathbb{E}\left[Z_{1,res}\right]$
according to Equation (A.35), implying that Equation (A.25) is valid.
Based on Equation (A.49), we have that
$\mathbb{E}\left[\tilde{Z}_{2,res}\right]$ $=\mathbb{E}\left[\tilde{U}_{2}\right]=\mathbb{E}\left[Z_{2,res}\right]$
according to Equation (A.36), implying that Equation (A.26) is valid.
Based on Equation (A.50), we have that
$Var\left(\tilde{Y}_{res}\right)$ $=\tilde{\beta}^{2}+Var\left(\tilde{\epsilon}_{res}\right)$
$=\tilde{\beta}^{2}+\left[Var\left(Y_{res}\right)-\tilde{\beta}^{2}\right]$
by Equation (A.38)
$=Var\left(Y_{res}\right),$
implying that Equation (A.27) is valid.
Based on Equation (A.51), we have that
$Var\left(\tilde{Z}_{1,res}\right)$
$\hskip 20.0pt=1+Var\left(\tilde{U}_{1}\right)+2\cdot Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{1}\right)$
$\hskip 20.0pt=1+\left[1+Var\left(Z_{1,res}\right)-2\cdot\dfrac{Cov\left(Z_{1,res},Y_{res}\right)}{\tilde{\beta}}\right]+2\cdot\left[\dfrac{Cov\left(Z_{1,res},Y_{res}\right)}{\tilde{\beta}}-1\right]$
by Equations (A.39) and (A.44)
$\hskip 20.0pt=Var\left(Z_{1,res}\right),$
implying that Equation (A.28) is valid.
Based on Equation (A.52), we have that
$Var\left(\tilde{Z}_{2,res}\right)$
$\hskip 20.0pt=1+Var\left(\tilde{U}_{2}\right)+2\cdot Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{2}\right)$
$\hskip 20.0pt=1+\left[1+Var\left(Z_{2,res}\right)-2\cdot\dfrac{Cov\left(Z_{2,res},Y_{res}\right)}{\tilde{\beta}}\right]+2\cdot\left[\dfrac{Cov\left(Z_{2,res},Y_{res}\right)}{\tilde{\beta}}-1\right]$
by Equations (A.40) and (A.45)
$\hskip 20.0pt=Var\left(Z_{2,res}\right),$
implying that Equation (A.29) is valid.
Based on Equation (A.53), we have that
$Cov\left(\tilde{Z}_{1,res},\tilde{Y}_{res}\right)$
$\hskip 20.0pt=\tilde{\beta}+\tilde{\beta}\cdot Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{1}\right)$
$\hskip 20.0pt=\tilde{\beta}+\tilde{\beta}\cdot\left[\dfrac{Cov\left(Z_{1,res},Y_{res}\right)}{\tilde{\beta}}-1\right]$
by Equation (A.44)
$\hskip 20.0pt=Cov\left(Z_{1,res},Y_{res}\right),$
implying that Equation (A.30) is valid.
Based on Equation (A.54), we have that
$Cov\left(\tilde{Z}_{2,res},\tilde{Y}_{res}\right)$
$\hskip 20.0pt=\tilde{\beta}+\tilde{\beta}\cdot Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{2}\right)$
$\hskip 20.0pt=\tilde{\beta}+\tilde{\beta}\cdot\left[\dfrac{Cov\left(Z_{2,res},Y_{res}\right)}{\tilde{\beta}}-1\right]$
by Equation (A.45)
$\hskip 20.0pt=Cov\left(Z_{2,res},Y_{res}\right),$
implying that Equation (A.31) is valid.
Based on Equation (A.55), we have that
$Cov\left(\tilde{Z}_{1,res},\tilde{Z}_{2,res}\right)$
$\hskip 20.0pt=1+Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{1}\right)+Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{2}\right)+Cov\left(\tilde{U}_{1},\tilde{U}_{2}\right)$
$\hskip 20.0pt=1+\left[\dfrac{Cov\left(Z_{1,res},Y_{res}\right)}{\tilde{\beta}}-1\right]+\left[\dfrac{Cov\left(Z_{2,res},Y_{res}\right)}{\tilde{\beta}}-1\right]$
[…]
by Equations (A.44)-(A.46)
$=Cov\left(Z_{1,res},Z_{2,res}\right),$
implying that Equation (A.32) is valid.
This concludes the proof for the case $\tilde{\beta}>0$ because we have constructed a distribution for the candidate latent variables that satisfies the model restrictions and the data restrictions.
Part 2: $\tilde{\beta}=0$
First, we note that $\tilde{\beta}=0$ implies that $LB=0$ because $LB\geq 0$ by Assumption 6. Consequently, we have that
$Cov\left(Z_{1,res},Y_{res}\right)=0$ (A.56)
and
$Cov\left(Z_{2,res},Y_{res}\right)=0.$ (A.57)
Now, we need to create a distribution for the candidate latent variables $\left(\tilde{X}_{res}^{*},\tilde{\epsilon}_{res},\tilde{U}_{1},\tilde{U}_{2}\right)$. The candidate observable variables $\left(\tilde{Y}_{res},\tilde{Z}_{1,res},\tilde{Z}_{1,res}\right)$ are created based on Equations (A.15)-(A.17).
We impose
$\left(\tilde{X}_{res}^{*},\tilde{\epsilon}_{res},\tilde{U}_{1},\tilde{U}_{2}\right)\sim N\left(\boldsymbol{\mu},\boldsymbol{\Omega}\right),$
i.e., $\left(\tilde{X}_{res}^{*},\tilde{\epsilon}_{res},\tilde{U}_{1},\tilde{U}_{2}\right)$ are jointly normally distributed with mean $\boldsymbol{\mu}$ and covariance matrix $\boldsymbol{\Omega}$.
The vector $\boldsymbol{\mu}$ is given by
$\mathbb{E}\left[\tilde{X}_{res}^{*}\right]$ $\coloneqq 0$ (A.58)
$\mathbb{E}\left[\tilde{\epsilon}_{res}\right]$ $\coloneqq 0$ (A.59)
$\mathbb{E}\left[\tilde{U}_{1}\right]$ $\coloneqq\mathbb{E}\left[Z_{1,res}\right]$ (A.60)
$\mathbb{E}\left[\tilde{U}_{1}\right]$ $\coloneqq\mathbb{E}\left[Z_{2,res}\right]$ (A.61)
and the covariance matrix $\boldsymbol{\Omega}$ is given by
$Var\left(\tilde{X}_{res}^{*}\right)$ $\coloneqq 1$ (A.62)
$Var\left(\tilde{\epsilon}_{res}\right)$ $\coloneqq Var\left(Y_{res}\right)$ (A.63)
$Var\left(\tilde{U}_{1}\right)$ $\coloneqq 1+Var\left(Z_{1,res}\right)$ (A.64)
$Var\left(\tilde{U}_{2}\right)$ $\coloneqq 1+Var\left(Z_{2,res}\right)$ (A.65)
$Cov\left(\tilde{X}_{res}^{*},\tilde{\epsilon}_{res}\right)$ $\coloneqq 0$ (A.66)
$Cov\left(\tilde{U}_{1},\tilde{\epsilon}_{res}\right)$ $\coloneqq 0$ (A.67)
$Cov\left(\tilde{U}_{2},\tilde{\epsilon}_{res}\right)$ $\coloneqq 0$ (A.68)
$Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{1}\right)$ $\coloneqq-1$ (A.69)
$Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{2}\right)$ $\coloneqq-1$ (A.70)
$Cov\left(\tilde{U}_{1},\tilde{U}_{2}\right)$ $\coloneqq Cov\left(Z_{1,res},Z_{2,res}\right)+1.$ (A.71)
Observe that the above distribution is a valid normal distribution because $Var\left(\tilde{\epsilon}_{res}\right)\geq 0$, $Var\left(\tilde{U}_{1}\right)\geq 0$ and $Var\left(\tilde{U}_{2}\right)\geq 0$ by construction.
Now, we must show that the candidate latent variables satisfy Equations (A.18)-(A.23). Notice that Restrictions (A.18)-(A.22) are satisfied by construction. Observe also that Restriction (A.23) is satisfied due to Assumption 6.
Lastly, we must show that the data restrictions (Equations (A.24)-(A.32)) are valid. Since we have shown that candidate latent variables satisfy Equations (A.18)-(A.23), we can follow the same steps used to prove Lemma A.1 to show that
$\mathbb{E}\left[\tilde{Y}_{res}\right]$ $=0$ (A.72)
$\mathbb{E}\left[\tilde{Z}_{1,res}\right]$ $=\mathbb{E}\left[\tilde{U}_{1}\right]$ (A.73)
$\mathbb{E}\left[\tilde{Z}_{2,res}\right]$ $=\mathbb{E}\left[\tilde{U}_{2}\right]$ (A.74)
$Var\left(\tilde{Y}_{res}\right)$ $=\tilde{\beta}^{2}+Var\left(\tilde{\epsilon}_{res}\right)$ (A.75)
$Var\left(\tilde{Z}_{1,res}\right)$ $=1+Var\left(\tilde{U}_{1}\right)+2\cdot Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{1}\right)$ (A.76)
$Var\left(\tilde{Z}_{2,res}\right)$ $=1+Var\left(\tilde{U}_{2}\right)+2\cdot Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{2}\right)$ (A.77)
$Cov\left(\tilde{Z}_{1,res},\tilde{Y}_{res}\right)$ $=\tilde{\beta}+\tilde{\beta}\cdot Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{1}\right)$ (A.78)
$Cov\left(\tilde{Z}_{2,res},\tilde{Y}_{res}\right)$ $=\tilde{\beta}+\tilde{\beta}\cdot Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{2}\right)$ (A.79)
$Cov\left(\tilde{Z}_{1,res},\tilde{Z}_{2,res}\right)$ $=1+Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{1}\right)+Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{2}\right)+Cov\left(\tilde{U}_{1},\tilde{U}_{2}\right).$ (A.80)
Based on Equation (A.72), we have that
$\mathbb{E}\left[\tilde{Y}_{res}\right]$ $=0=\mathbb{E}\left[Y_{res}\right]$
according to Equation (A.1), implying that Equation (A.24) is valid.
Based on Equation (A.73), we have that
$\mathbb{E}\left[\tilde{Z}_{1,res}\right]$ $=\mathbb{E}\left[\tilde{U}_{1}\right]=\mathbb{E}\left[Z_{1,res}\right]$
according to Equation (A.60), implying that Equation (A.25) is valid.
Based on Equation (A.74), we have that
$\mathbb{E}\left[\tilde{Z}_{2,res}\right]$ $=\mathbb{E}\left[\tilde{U}_{2}\right]=\mathbb{E}\left[Z_{2,res}\right]$
according to Equation (A.61), implying that Equation (A.26) is valid.
Based on Equation (A.75), we have that
$Var\left(\tilde{Y}_{res}\right)$ $=Var\left(\tilde{\epsilon}_{res}\right)=Var\left(Y_{res}\right)$
by Equation (A.63), implying that Equation (A.27) is valid.
Based on Equation (A.76), we have that
$Var\left(\tilde{Z}_{1,res}\right)$
$\hskip 20.0pt=1+Var\left(\tilde{U}_{1}\right)+2\cdot Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{1}\right)$
$\hskip 20.0pt=1+\left[1+Var\left(Z_{1,res}\right)\right]+2\cdot\left[-1\right]$
by Equations (A.64) and (A.69)
$\hskip 20.0pt=Var\left(Z_{1,res}\right),$
implying that Equation (A.28) is valid.
Based on Equation (A.77), we have that
$Var\left(\tilde{Z}_{2,res}\right)$
$\hskip 20.0pt=1+Var\left(\tilde{U}_{2}\right)+2\cdot Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{2}\right)$
$\hskip 20.0pt=1+\left[1+Var\left(Z_{2,res}\right)\right]+2\cdot\left[-1\right]$
by Equations (A.65) and (A.70)
$\hskip 20.0pt=Var\left(Z_{2,res}\right),$
implying that Equation (A.29) is valid.
Based on Equation (A.78), we have that
$Cov\left(\tilde{Z}_{1,res},\tilde{Y}_{res}\right)$ $=0=Cov\left(Z_{1,res},Y_{res}\right)$
by Equation (A.56), implying that Equation (A.30) is valid.
Based on Equation (A.79), we have that
$Cov\left(\tilde{Z}_{2,res},\tilde{Y}_{res}\right)$ $=0=Cov\left(Z_{2,res},Y_{res}\right)$
by Equation (A.57), implying that Equation (A.31) is valid.
Based on Equation (A.55), we have that
$Cov\left(\tilde{Z}_{1,res},\tilde{Z}_{2,res}\right)$
$\hskip 20.0pt=1+Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{1}\right)+Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{2}\right)+Cov\left(\tilde{U}_{1},\tilde{U}_{2}\right)$
$\hskip 20.0pt=1+\left[-1\right]+\left[-1\right]+\left[1+Cov\left(Z_{1,res},Z_{2,res}\right)\right]$
by Equations (A.69)-(A.71)
$=Cov\left(Z_{1,res},Z_{2,res}\right),$
implying that Equation (A.32) is valid.
This concludes the proof for the case $\tilde{\beta}=0$ because we have constructed a distribution for the candidate latent variables that satisfies the model restrictions and the data restrictions.
Combining Parts 1 and 2 of Appendix A.1.4, we can conclude that, if $UB\geq LB$, then the interval $\left[LB,UB\right]$ exhausts all the information contained in the first two moments of the data distribution.
Appendix B Simple Examples
In this appendix, we propose three simple data-generating processes where alternative methodologies perform poorly when compared to our partial identification strategy. Our first example (Appendix B.1) shows that the reduced-form coefficient (Equation (7)) may be strictly less than our lower bound, while our second example (Appendix B.2) shows that the reduced-form coefficient may be strictly greater than our upper bound. Lastly, our third example (Appendix B.3) shows that the estimand proposed by Lubotsky and Wittenberg (2006) may present a larger bias than our lower bound.
Example 1: $b_{1}<LB$
In this section, we create a class of data-generating processes such that
$b_{1}<LB,$
where $b_{1}$ is defined in Equation (7).
To do so, assume that $\beta>0$, $c_{1}>0$ and
$U_{1}=a_{1}+c_{1}\cdot X_{res}^{*},$ (B.1)
implying that
$Z_{1,res}=a_{1}+\left(1+c_{1}\right)\cdot X_{res}^{*}$ (B.2)
according to Equation (3). In this case, we have that
$b_{1}$ $=\dfrac{Cov\left(Z_{1,res},Y_{res}\right)}{Var\left(Z_{1,res}\right)}$
by definition
$=\dfrac{Cov\left(a_{1}+\left(1+c_{1}\right)\cdot X_{res}^{*},\hskip 10.0pt\beta\cdot X_{res}^{*}+\epsilon_{res}\right)}{Var\left(a_{1}+\left(1+c_{1}\right)\cdot X_{res}^{*}\right)}$
according to Equations (2) and (B.2)
$=\dfrac{\beta\cdot\left(1+c_{1}\right)}{\left(1+c_{1}\right)^{2}}$ (B.3)
[…]
Moreover, our lower bound satisfies:
$LB$ $\geq 2\cdot\dfrac{Cov\left(Z_{1,res},Y_{res}\right)}{1+Var\left(Z_{1,res}\right)}$
according to Equation (4)
$=2\cdot\dfrac{\beta\cdot\left(1+c_{1}\right)}{1+\left(1+c_{1}\right)^{2}}$ (B.4)
according to the same steps used to derived Equation (B.3).
Now, note that
[…]
Intuitively, the last result implies that, if the covariance between the true unobservable variable $X_{res}^{*}$ and its proxy $Z_{1,res}$ is sufficiently large, then the reduced-form coefficient will be more biased than our lower bound.
Therefore, we can conclude that there exist simple data-generating processes where our partial identification strategy’s performance compares positively against the reduce-form approach.
Example 2: $b_{1}>UB$
In this section, we create a class of data-generating processes such that
$b_{1}>UB,$
where $b_{1}$ is defined in Equation (7).
To do so, assume that $\beta=UB\coloneqq\sqrt{Var\left(Y\right)}$, $\epsilon_{res}=0$ and
$\left[\begin{matrix}X_{res}^{*}\\ U_{1}\end{matrix}\right]\sim N\left(\left[\begin{matrix}0\\ 0\end{matrix}\right],\left[\begin{matrix}1&Cov\left(X_{res}^{*},U_{1}\right)\\ Cov\left(X_{res}^{*},U_{1}\right)&Var\left(U_{1}\right)\end{matrix}\right]\right),$
where $Var\left(U_{1}\right)\in\left(0,1\right)$ and $Cov\left(X_{res}^{*},U_{1}\right)\in\left(-\dfrac{1+Var\left(U_{1}\right)}{2},-Var\left(U_{1}\right)\right)$.
In this case, we have that
$b_{1}$ $=\dfrac{Cov\left(Z_{1,res},Y_{res}\right)}{Var\left(Z_{1,res}\right)}$
by definition
$=\dfrac{Cov\left(X_{res}^{*}+U_{1},\beta\cdot X_{res}^{*}+\epsilon_{res}\right)}{Var\left(X_{res}^{*}+U_{1}\right)}$
by Equations (2) and (3)
$=\dfrac{Cov\left(X_{res}^{*}+U_{1},UB\cdot X_{res}^{*}\right)}{Var\left(X_{res}^{*}+U_{1}\right)}$
by assumption
$=\dfrac{UB+UB\cdot Cov\left(X_{res}^{*},U_{1}\right)}{1+Var\left(U_{1}\right)+2\cdot Cov\left(X_{res}^{*},U_{1}\right)}$
by Assumption 2
$=UB\cdot\left[\dfrac{1+Cov\left(X_{res}^{*},U_{1}\right)}{1+Var\left(U_{1}\right)+2\cdot Cov\left(X_{res}^{*},U_{1}\right)}\right]$
$\hskip 20.0pt\text{where }\dfrac{1+Cov\left(X_{res}^{*},U_{1}\right)}{1+Var\left(U_{1}\right)+2\cdot Cov\left(X_{res}^{*},U_{1}\right)}>0$
$\hskip 20.0pt\text{because }Cov\left(X_{res}^{*},U_{1}\right)>-\dfrac{1+Var\left(U_{1}\right)}{2}>-1$
$>UB\cdot\left[\dfrac{1+Cov\left(X_{res}^{*},U_{1}\right)}{1-Cov\left(X_{res}^{*},U_{1}\right)+2\cdot Cov\left(X_{res}^{*},U_{1}\right)}\right]$
$\hskip 20.0pt\text{because }Cov\left(X_{res}^{*},U_{1}\right)<-Var\left(U_{1}\right)$
$=UB=\beta.$
Intuitively, the last result implies that, if the covariance between the true unobservable variable $X_{res}^{*}$ and the measurement error $U_{1}$ is neither too large nor too small in comparison to the variance of the measurement error, then the reduce-form coefficient will be more biased than our upper bound.
Therefore, we can conclude that there exist simple data-generating processes where our partial identification strategy’s performance compares positively against the reduce-form approach.
Example 3: Comparing our method against the one proposed by Lubotsky and Wittenberg (2006)
In this section, we compare our method to the closest alternative tool in the literature: the bias minimizing linear regression estimand proposed by Lubotsky and Wittenberg (2006). We start by explaining the method developed by these authors and, then, describe a data-generating process where our lower bound is closer to the target parameter than the estimand proposed by Lubotsky and Wittenberg (2006).
Similarly to us, Lubotsky and Wittenberg (2006) observe two proxies for an unobservable variable of interest, $X^{*}$, and are interested in the coefficient of the linear regression model described by Equation (2). Differently from us, they do not restrict the covariance between the error terms, i.e., they do not impose Assumption 5. For this reason, we compare their procedure against the method described in Appendix C, which does not impose this assumption either. In the current appendix, we impose Assumptions 1, 2, 3, 4 and a weaker version of Assumption 6. Using these assumptions, Proposition C.1 shows that
$\beta\geq LB_{C}\coloneqq\max\left\{\begin{matrix}2\cdot\dfrac{Cov\left(Z_{1,res},Y_{res}\right)}{1+Var\left(Z_{1,res}\right)},\hskip 10.0pt2\cdot\dfrac{Cov\left(Z_{2,res},Y_{res}\right)}{1+Var\left(Z_{2,res}\right)}\end{matrix}\right\}.$
Lubotsky and Wittenberg (2006) impose Assumptions 1, 4 and an additional restriction on the covariance between the error terms and the true unobservable variable of interest. Specifically, they impose the following condition.
Assumption B.1 (Individually Classical Measurement Errors)
Each measurement error term is individually classic, i.e., $Cov\left(X^{*}_{res},U_{k}\right)=0$ for any $k\in\left\{1,2\right\}$.
Note that this assumption does not impose that the measurement error terms are jointly classical since $Cov\left(U_{1},U_{2}\right)$ might be different from zero. As a consequence, using $Z_{2,res}$ as an instrument for $Z_{1,res}$ does not identify the target parameter.
Moreover, observe that Lubotsky and Wittenberg (2006) do not restrict neither the variance of the true unobservable variable nor the sign of the true coefficient. However, to simplify the exposition, we will impose Assumptions 2 and 3 in this appendix.
To overcome the identification challenges posed by measurement error, Lubotsky and Wittenberg (2006) propose to use convex combinations of the two proxies in a linear regression model. Specifically, they define
$Z\left(\delta\right)\coloneqq\delta\cdot Z_{1,res}+\left(1-\delta\right)\cdot Z_{2,res},$
where $\delta\in\left[0,1\right]$, and analyze
$Y_{res}=b\left(\delta\right)\cdot X^{*}_{res}+V_{\delta},$
where $b\left(\delta\right)\coloneqq\frac{Cov\left(Z\left(\delta\right),Y_{res}\right)}{Var\left(Z\left(\delta\right)\right)}$ and $V_{\delta}$ is the residual. Due to individually classic measurement error terms and Assumption 3, we have that
$\beta\geq b\left(\delta\right)$
for any $\delta\in\left[0,1\right]$.
To minimize bias, Lubotsky and Wittenberg (2006) propose to focus on the convex combination of proxy variables that generates the OLS coefficient that is as close as possible to the target parameter. In other words, they want to identify
$b^{*}\coloneqq\operatorname*{\arg\!\max}\limits_{\delta\in\left[0,1\right]}b\left(\delta\right).$
Interestingly, they show that one can identify $b^{*}$ according to
$b^{*}=b_{1}+b_{2},$
where $b_{1}$ and $b_{2}$ are defined as the OLS coefficients of the following linear regression model:
$Y_{res}=b_{1}\cdot Z_{1,res}+b_{2}\cdot Z_{2,res}+V.$
Under Assumptions 1, 2, 4 and B.1, Lubotsky and Wittenberg (2006, Equation (4)) shows that
$b^{*}=\beta\cdot\left[\frac{Var\left(U_{1}\right)+Var\left(U_{2}\right)-2\cdot Cov\left(U_{1},U_{2}\right)}{Var\left(U_{1}\right)+Var\left(U_{2}\right)-3\cdot Cov\left(U_{1},U_{2}\right)+Var\left(U_{1}\right)\cdot Var\left(U_{2}\right)}\right]$ (B.5)
Moreover, under Assumptions 1, 2, 3, 4, 6 and B.1, our lower bound satisfies
$LB_{C}$ $\geq 2\cdot\dfrac{Cov\left(Z_{1,res},Y_{res}\right)}{1+Var\left(Z_{1,res}\right)}$
according to Proposition C.1
$=\beta\cdot\left[\frac{2}{2+Var\left(U_{1}\right)}\right]$ (B.6)
according to Lemma C.1.
Now, we show that there exist data-generating processes that satisfy Assumptions 1, 2, 3, 4, 6 and B.1 such that $\beta\geq LB_{C}>b^{*}$.
Assume that $Var\left(U_{1}\right)=1$, $Var\left(U_{2}\right)>1$ and $Var\left(U_{2}\right)\geq\frac{3\cdot Cov\left(U_{1},U_{2}\right)-1}{2}$. Notice that Equation (B.6) implies that
$LB_{C}\geq\frac{2}{3}\cdot\beta$ (B.7)
and Equation (B.5) implies that
$b^{*}=\beta\cdot\left[\frac{1+Var\left(U_{2}\right)-2\cdot Cov\left(U_{1},U_{2}\right)}{1+2\cdot Var\left(U_{2}\right)-3\cdot Cov\left(U_{1},U_{2}\right)}\right].$ (B.8)
Observe also that
$\begin{array}[]{rrcl}&Var\left(U_{2}\right)&>&1\\ \Leftrightarrow&2+4\cdot Var\left(U_{2}\right)-6\cdot Cov\left(U_{1},U_{2}\right)&>&3+3\cdot Var\left(U_{2}\right)-6\cdot Cov\left(U_{1},U_{2}\right)\\ \Leftrightarrow&\frac{2}{3}\cdot\left[1+2\cdot Var\left(U_{2}\right)-3\cdot Cov\left(U_{1},U_{2}\right)\right]&>&1+1\cdot Var\left(U_{2}\right)-2\cdot Cov\left(U_{1},U_{2}\right)\\ \Leftrightarrow&\frac{2}{3}&>&\frac{1+1\cdot Var\left(U_{2}\right)-2\cdot Cov\left(U_{1},U_{2}\right)}{1+2\cdot Var\left(U_{2}\right)-3\cdot Cov\left(U_{1},U_{2}\right)}.\end{array}$
because $Var\left(U_{2}\right)\geq\frac{3\cdot Cov\left(U_{1},U_{2}\right)-1}{2}$.
Combining the last equation with Equations (B.7)-(B.8) and Assumption 3, we find that $\beta\geq LB_{C}>b^{*}$. In plain English, in this data-generating process, our lower bound presents a smaller bias than the estimand proposed by Lubotsky and Wittenberg (2006).
Intuitively, we achieve a smaller bias than $b^{*}$ because our lower bound does not restrict itself to convex combinations of the proxy variables. For example, the object $l_{1}\coloneqq 2\cdot\dfrac{Cov\left(Z_{1,res},Y_{res}\right)}{1+Var\left(Z_{1,res}\right)}$ can be seen as the coefficient of the following linear regression model:
$Y_{res}=l_{1}\cdot\tilde{Z}+V_{1},$
where $\tilde{Z}\coloneqq\frac{Z_{1,res}+\tilde{V}}{2}$ and $\tilde{V}$ is a random variable satisfying $Cov\left(\tilde{V},Y\right)=0$, $Cov\left(Z_{1,res},\tilde{V}\right)=0$ and $Var\left(\tilde{V}\right)=1$. In other words, $\tilde{Z}$ is the average of $Z_{1,res}$ and a noise term that has the same variance of $X_{res}^{*}$.
Therefore, we can conclude that there exist simple data-generating processes that satisfy Assumptions 1, 2, 3, 4, 6 and B.1 and create a scenario where our partial identification strategy’s performance compares positively against the procedure recommended by Lubotsky and Wittenberg (2006).
Appendix C Unrestricted Covariance between the Measurement Error Terms
In this appendix, we propose a partial identification strategy that does not restrict the covariance between the proxies’s measurement error terms $\left(U_{1},U_{2}\right)$. To do so, we keep Assumptions 1, 2, 3 and 4, completely remove Assumption 5, and modify Assumption 6. For convenience, we list all the necessary assumptions here.
Assumption C.1 (Exogeneity)
After conditioning on the covariate vector $W$, the true unobserved measure $X^{*}$ is exogenous, i.e., $Cov\left(X_{res}^{*},\epsilon_{res}\right)=0$.
Assumption C.2 (Scale Normalization)
The scale of the true unobserved measure is normalized to have unit variance, i.e., $Var\left(X_{res}^{*}\right)=1$.
Assumption C.3 (Known Direction of the True Effect)
The direction of the effect of the true unobservable variable of interest on the outcome variable is known, i.e., $\beta\geq 0$.
Assumption C.4 (Exogeneity of the Measurement Errors)
The measurement error variables are exogenous, i.e., $Cov\left(U_{1},\epsilon_{res}\right)=0$ and $Cov\left(U_{2},\epsilon_{res}\right)=0$
Assumption C.5 (Well-behaved Data Distributions)
The variance of the outcome variable is strictly positive and the covariances between each proxy and the outcome variable are positive, i.e., $Var\left(Y_{res}\right)>0$, $Cov\left(Z_{1,res},Y_{res}\right)\geq 0$ and $Cov\left(Z_{2,res},Y_{res}\right)\geq 0.$
We now state our main result: partial identification of $\beta$ (Equation (2)) using two proxies (Equation (3)) with nonclassical measurement error and unrestricted covariance between the measurement error terms.
Proposition C.1
Under Assumptions C.1-C.5, we have that:
$\beta\in\left\{\begin{array}[]{cl}\left\{\emptyset\right\}&\text{ if }UB<LB\\ \left[LB,UB\right]&\text{ if }UB\geq LB\end{array}\right$., (C.1)
where
$LB_{C}\coloneqq\max\left\{\begin{matrix}2\cdot\dfrac{Cov\left(Z_{1,res},Y_{res}\right)}{1+Var\left(Z_{1,res}\right)},\hskip 10.0pt2\cdot\dfrac{Cov\left(Z_{2,res},Y_{res}\right)}{1+Var\left(Z_{2,res}\right)}\end{matrix}\right\}$ (C.2)
and
$UB\coloneqq\sqrt{Var\left(Y_{res}\right)},$ (C.3)
and note that $LB_{C}\geq 0$ according to Assumption C.5.
Moreover, the bounds in Equation (C.1) exhaust all the information contained in the first two moments of the observable variables $\left(Y_{res},Z_{1,res},Z_{2,res}\right)$.
Proof.
See Appendix C.1. ∎
The only difference between the bounds in Proposition C.1 and the bounds in Proposition 1 is the lower bound. Since the term $\dfrac{Cov\left(Z_{1,res},Y_{res}\right)+Cov\left(Z_{2,res},Y_{res}\right)}{1+Cov\left(Z_{1,res},Z_{2,res}\right)}$ captures the identifying power of Assumption 5, it is not present in $LB_{C}$. Consequently, we have that that $LB_{C}\leq LB$, i.e., the bounds in Proposition C.1 are weakly wider than the bounds in Proposition 1.
Proof of Proposition C.1
We break the proof of Proposition C.1 in 4 parts:
  1. 1.
    Section C.1.1 derives the restrictions imposed by the first two moments of the observable data on the underlying data-generating process.
  2. 2.
    Section C.1.2 shows that $LB_{C}\leq\beta\leq UB$ if $UB\geq LB_{C}$.
  3. 3.
    Section C.1.3 shows that, if $LB_{C}>UB$, then there is no data-generating process that satisfies Assumptions C.1-C.5 and the restrictions imposed by the first two moments of the data distribution.
  4. 4.
    Section C.1.4 defines the meaning of exhausting all the information contained in the first two moments of the data distribution and shows that, if $UB\geq LB_{C}$, then the interval $\left[LB_{C},UB\right]$ exhausts all the information contained in the first two moments of the data distribution.
C.1.1 Data Restrictions
In this section, we derive the restrictions imposed by the first two moments of the observable data distribution $\left(Y_{res},Z_{1,res},Z_{2,res}\right)$ on the distribution of the latent variables $\left(X_{res}^{*},\epsilon_{res},U_{1},U_{2}\right)$. These restrictions are the same restrictions described in Appendix A.1.1. For this reason, we re-state them here for convenience, but do not provide a proof.
Lemma C.1 (Data Restrictions)
Under Assumptions C.1, C.2 and C.4, we have that
$\mathbb{E}\left[Y_{res}\right]$ $=0$ (C.4)
$\mathbb{E}\left[Z_{1,res}\right]$ $=\mathbb{E}\left[U_{1}\right]$ (C.5)
$\mathbb{E}\left[Z_{2,res}\right]$ $=\mathbb{E}\left[U_{2}\right]$ (C.6)
$Var\left(Y_{res}\right)$ $=\beta^{2}+Var\left(\epsilon_{res}\right)$ (C.7)
$Var\left(Z_{1,res}\right)$ $=1+Var\left(U_{1}\right)+2\cdot Cov\left(X_{res}^{*},U_{1}\right)$ (C.8)
$Var\left(Z_{2,res}\right)$ $=1+Var\left(U_{2}\right)+2\cdot Cov\left(X_{res}^{*},U_{2}\right)$ (C.9)
$Cov\left(Z_{1,res},Y_{res}\right)$ $=\beta+\beta\cdot Cov\left(X_{res}^{*},U_{1}\right)$ (C.10)
$Cov\left(Z_{2,res},Y_{res}\right)$ $=\beta+\beta\cdot Cov\left(X_{res}^{*},U_{2}\right)$ (C.11)
$Cov\left(Z_{1,res},Z_{2,res}\right)$ $=1+Cov\left(X_{res}^{*},U_{1}\right)+Cov\left(X_{res}^{*},U_{2}\right)+Cov\left(U_{1},U_{2}\right).$ (C.12)
C.1.2 $LB_{C}\leq\beta\leq UB$ if $UB\geq LB_{C}$
In this section, we consider the case when $UB\geq LB_{C}$ and prove that (i) $\beta\leq UB$, (ii) $\beta\geq LB_{C}$ when $\beta=0$, and (iii) $\beta\geq LB_{C}$ when $\beta>0$. In this proof, we do not impose that $LB_{C}\geq 0$.
Part 1: $\beta\leq UB$
Rearranging Equation (C.7), we have that
$\beta^{2}$ $=Var\left(Y_{res}\right)-Var\left(\epsilon_{res}\right)$
$\leq Var\left(Y_{res}\right)$
because $Var\left(\epsilon_{res}\right)\geq 0$, implying that
$\beta\leq\sqrt{Var\left(Y_{res}\right)}=UB$
by Assumption C.3.
Part 2: $\beta\geq LB_{C}$ when $\beta=0$
By combining Equations (C.10) and (C.11) with $\beta=0$, we find that
$Cov\left(Z_{1,res},Y_{res}\right)=0$
and
$Cov\left(Z_{2,res},Y_{res}\right)=0,$
implying that
$LB_{C}=0,$
according to Equation (C.2), and that
$\beta=0=LB_{C}.$
Part 3: $\beta\geq LB_{C}$ when $\beta>0$
Having in mind that $\beta>0$ and rearranging Equations (C.10) and (C.11), we have that
$Cov\left(X_{res}^{*},U_{1}\right)=\dfrac{Cov\left(Z_{1,res},Y_{res}\right)}{\beta}-1$ (C.13)
and
$Cov\left(X_{res}^{*},U_{2}\right)=\dfrac{Cov\left(Z_{2,res},Y_{res}\right)}{\beta}-1.$ (C.14)
Rearranging Equation (C.8), we find that
$Var\left(U_{1}\right)$ $=Var\left(Z_{1,res}\right)-1-2\cdot Cov\left(X_{res}^{*},U_{1}\right)$
$=Var\left(Z_{1,res}\right)-1-2\cdot\left[\dfrac{Cov\left(Z_{1,res},Y_{res}\right)}{\beta}-1\right]$
according to Equation (C.13)
$=1+Var\left(Z_{1,res}\right)-2\cdot\dfrac{Cov\left(Z_{1,res},Y_{res}\right)}{\beta}$
Since $Var\left(U_{1}\right)\geq 0$ and $\beta>0$, the last equation implies that
$\begin{array}[]{crcl}&1+Var\left(Z_{1,res}\right)-2\cdot\dfrac{Cov\left(Z_{1,res},Y_{res}\right)}{\beta}&\geq&0\\ \Leftrightarrow&1+Var\left(Z_{1,res}\right)&\geq&2\cdot\dfrac{Cov\left(Z_{1,res},Y_{res}\right)}{\beta}\\ \Leftrightarrow&\beta\cdot\left(1+Var\left(Z_{1,res}\right)\right)&\geq&2\cdot Cov\left(Z_{1,res},Y_{res}\right),\end{array}$
implying that
$\beta\geq 2\cdot\dfrac{Cov\left(Z_{1,res},Y_{res}\right)}{1+Var\left(Z_{1,res}\right)}.$ (C.15)
Combining Equations (C.9) and (C.14) and using the same steps used to derive Equation (C.15), we have that
$\beta\geq 2\cdot\dfrac{Cov\left(Z_{2,res},Y_{res}\right)}{1+Var\left(Z_{2,res}\right)}.$ (C.16)
Combining Equations (C.2), (C.15) and (C.16), we have that
$\beta\geq LB_{C}.$
C.1.3 No Valid Data-Generating Process if $LB_{C}>UB$
In this section, we show that, when $LB_{C}>UB$, there is no combination of $\beta\geq 0$ and distribution of latent variables $\left(X_{res}^{*},\epsilon_{res},U_{1},U_{2}\right)$ that satisfies Assumptions C.1-C.5 and the data restrictions described in Lemma C.1. We prove this result by contradiction.
Assume that $LB_{C}>UB$ and that there exist parameter $\beta\geq 0$ and variables $\left(X_{res}^{*},\epsilon_{res},U_{1},U_{2}\right)$ such that Assumptions C.1-C.5 and Equations (C.4)-(C.12) are valid.
Following the same steps described in Parts 2 and 3 in Section C.1.2, we must have that $\beta\geq LB_{C}$. Rearranging Equation (C.7), we have that
$Var\left(\epsilon_{res}\right)$ $=Var\left(Y_{res}\right)-\beta^{2}$
$\leq Var\left(Y_{res}\right)-\left(LB_{C}\right)^{2}$
[…]
$=\left(UB\right)^{2}-\left(LB_{C}\right)^{2}$
according to Equation (C.3)
$<0\hskip 10.0pt\text{because }LB_{C}>UB,$
which is a contradiction.
C.1.4 $\left[LB_{C},UB\right]$ exhausts all the information contained in the first two moments of the data distribution if $UB\geq LB_{C}$
First, we define the meaning of exhausting all the information contained in the first two moments of the data distribution. Second, we show that, if $UB\geq LB_{C}$, then the interval $\left[LB_{C},UB\right]$ exhausts all the information contained in the first two moments of the data distribution.
Definition C.1
A set $\mathcal{B}\subseteq\mathbb{R}$ exhausts all the information contained in the first two moments of the data distribution if, for any candidate target parameter $\tilde{\beta}\in\mathcal{B}$, there exist candidate latent variables $\left(\tilde{X}_{res}^{*},\tilde{\epsilon}_{res},\tilde{U}_{1},\tilde{U}_{2}\right)$ and candidate observable variables $\left(\tilde{Y}_{res},\tilde{Z}_{1,res},\tilde{Z}_{1,res}\right)$ such that
  1. 1.
    the candidate observable variables are induced by the candidate latent variables according to the model restrictions imposed by Equations (2) and (3), i.e.,
    $\tilde{Y}_{res}$ $\coloneqq\tilde{\beta}\cdot\tilde{X}_{res}^{*}+\tilde{\epsilon}_{res}$ (C.17)
    $\tilde{Z}_{1,res}$ $\coloneqq\tilde{X}_{res}^{*}+\tilde{U}_{1}$ (C.18)
    $\tilde{Z}_{2,res}$ $\coloneqq\tilde{X}_{res}^{*}+\tilde{U}_{2};$ (C.19)
  2. 2.
    the candidate latent variables and the candidate target parameter satisfy the model restrictions imposed by Assumptions C.1-C.4, i.e.,
    $Cov\left(\tilde{X}_{res}^{*},\tilde{\epsilon}_{res}\right)$ $=0$ (C.20)
    $Var\left(\tilde{X}_{res}^{*}\right)$ $=1$ (C.21)
    $\tilde{\beta}$ $\geq 0$ (C.22)
    $Cov\left(\tilde{U}_{1},\tilde{\epsilon}_{res}\right)$ $=0$ (C.23)
    $Cov\left(\tilde{U}_{2},\tilde{\epsilon}_{res}\right)$ $=0$ (C.24)
  3. 3.
    the first two moments of the candidate observable variables are equal to the first two moments of the true observable variables, i.e.,
    $\mathbb{E}\left[\tilde{Y}_{res}\right]$ $=\mathbb{E}\left[Y_{res}\right]$ (C.25)
    $\mathbb{E}\left[\tilde{Z}_{1,res}\right]$ $=\mathbb{E}\left[Z_{1,res}\right]$ (C.26)
    $\mathbb{E}\left[\tilde{Z}_{2,res}\right]$ $=\mathbb{E}\left[Z_{2,res}\right]$ (C.27)
    $Var\left(\tilde{Y}_{res}\right)$ $=Var\left(Y_{res}\right)$ (C.28)
    $Var\left(\tilde{Z}_{1,res}\right)$ $=Var\left(Z_{1,res}\right)$ (C.29)
    $Var\left({Z}_{2,res}\right)$ $=Var\left(Z_{2,res}\right)$ (C.30)
    $Cov\left(\tilde{Z}_{1,res},\tilde{Y}_{res}\right)$ $=Cov\left(Z_{1,res},Y_{res}\right)$ (C.31)
    $Cov\left(\tilde{Z}_{2,res},\tilde{Y}_{res}\right)$ $=Cov\left(Z_{2,res},Y_{res}\right)$ (C.32)
    $Cov\left(\tilde{Z}_{1,res},\tilde{Z}_{2,res}\right)$ $=Cov\left(Z_{1,res},Z_{2,res}\right).$ (C.33)
Second, we show that, if $UB\geq LB_{C}\geq 0$, then the interval $\left[LB_{C},UB\right]$ exhausts all the information contained in the first two moments of the data distribution.
Assume that $UB\geq LB_{C}$ and fix $\tilde{\beta}\in\left[LB_{C},UB\right]$ arbitrarily. We consider two cases: (i) $\tilde{\beta}>0$, and (ii) $\tilde{\beta}=0$. In both cases, the proof is by construction.
Part 1: $\tilde{\beta}>0$
We need to create a distribution for the candidate latent variables $\left(\tilde{X}_{res}^{*},\tilde{\epsilon}_{res},\tilde{U}_{1},\tilde{U}_{2}\right)$. The candidate observable variables $\left(\tilde{Y}_{res},\tilde{Z}_{1,res},\tilde{Z}_{1,res}\right)$ are created based on Equations (C.17)-(C.19).
We impose
$\left(\tilde{X}_{res}^{*},\tilde{\epsilon}_{res},\tilde{U}_{1},\tilde{U}_{2}\right)\sim N\left(\boldsymbol{\mu},\boldsymbol{\Omega}\right),$
i.e., $\left(\tilde{X}_{res}^{*},\tilde{\epsilon}_{res},\tilde{U}_{1},\tilde{U}_{2}\right)$ are jointly normally distributed with mean $\boldsymbol{\mu}$ and covariance matrix $\boldsymbol{\Omega}$.
The vector $\boldsymbol{\mu}$ is given by
$\mathbb{E}\left[\tilde{X}_{res}^{*}\right]$ $\coloneqq 0$ (C.34)
$\mathbb{E}\left[\tilde{\epsilon}_{res}\right]$ $\coloneqq 0$ (C.35)
$\mathbb{E}\left[\tilde{U}_{1}\right]$ $\coloneqq\mathbb{E}\left[Z_{1,res}\right]$ (C.36)
$\mathbb{E}\left[\tilde{U}_{1}\right]$ $\coloneqq\mathbb{E}\left[Z_{2,res}\right]$ (C.37)
and the covariance matrix $\boldsymbol{\Omega}$ is given by
$Var\left(\tilde{X}_{res}^{*}\right)$ $\coloneqq 1$ (C.38)
$Var\left(\tilde{\epsilon}_{res}\right)$ $\coloneqq Var\left(Y_{res}\right)-\tilde{\beta}^{2}$ (C.39)
$Var\left(\tilde{U}_{1}\right)$ $\coloneqq 1+Var\left(Z_{1,res}\right)-2\cdot\dfrac{Cov\left(Z_{1,res},Y_{res}\right)}{\tilde{\beta}}$ (C.40)
$Var\left(\tilde{U}_{2}\right)$ $\coloneqq 1+Var\left(Z_{2,res}\right)-2\cdot\dfrac{Cov\left(Z_{2,res},Y_{res}\right)}{\tilde{\beta}}$ (C.41)
$Cov\left(\tilde{X}_{res}^{*},\tilde{\epsilon}_{res}\right)$ $\coloneqq 0$ (C.42)
$Cov\left(\tilde{U}_{1},\tilde{\epsilon}_{res}\right)$ $\coloneqq 0$ (C.43)
$Cov\left(\tilde{U}_{2},\tilde{\epsilon}_{res}\right)$ $\coloneqq 0$ (C.44)
$Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{1}\right)$ $\coloneqq\dfrac{Cov\left(Z_{1,res},Y_{res}\right)}{\tilde{\beta}}-1$ (C.45)
$Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{2}\right)$ $\coloneqq\dfrac{Cov\left(Z_{2,res},Y_{res}\right)}{\tilde{\beta}}-1$ (C.46)
$Cov\left(\tilde{U}_{1},\tilde{U}_{2}\right)$ $\coloneqq 1+Cov\left(Z_{1,res},Z_{2,res}\right)-\left[\dfrac{Cov\left(Z_{1,res},Y_{res}\right)+Cov\left(Z_{2,res},Y_{res}\right)}{\tilde{\beta}}\right$] (C.47)
First, we must show that the above distribution is a valid normal distribution, i.e., we must show that $Var\left(\tilde{\epsilon}_{res}\right)\geq 0$, $Var\left(\tilde{U}_{1}\right)\geq 0$ and $Var\left(\tilde{U}_{2}\right)\geq 0$.
Note that
$Var\left(\tilde{\epsilon}_{res}\right)$ $\coloneqq Var\left(Y_{res}\right)-\tilde{\beta}^{2}$
according to Equation (C.39)
$\geq Var\left(Y_{res}\right)-\left(UB\right)^{2}$
$\hskip 20.0pt\text{because }\tilde{\beta}\leq UB$
$=Var\left(Y_{res}\right)-\left(\sqrt{Var\left(Y_{res}\right)}\right)^{2}$
according to Equation (C.3)
$=0.$
Now, we show that $Var\left(\tilde{U}_{1}\right)\geq 0$. To do so, we analyze two cases. If $Cov\left(Z_{1,res},Y_{res}\right)=0$, then $Var\left(\tilde{U}_{1}\right)=1+Var\left(Z_{1,res}\right)>0$ according to Equation (C.40). If $Cov\left(Z_{1,res},Y_{res}\right)>0$, then
$Var\left(\tilde{U}_{1}\right)$ $=1+Var\left(Z_{1,res}\right)-2\cdot\dfrac{Cov\left(Z_{1,res},Y_{res}\right)}{\tilde{\beta}}$
according to Equation (C.40)
$\geq 1+Var\left(Z_{1,res}\right)-2\cdot\dfrac{Cov\left(Z_{1,res},Y_{res}\right)}{\nicefrac{{2\cdot Cov\left(Z_{1,res},Y_{res}\right)}}{{\left(1+Var\left(Z_{1,res}\right)\right)}}}$
[…]
$=0.$
We can show that $Var\left(\tilde{U}_{2}\right)\geq 0$ using Equation (C.41) and the steps above.
Hence, we can conclude that the proposed distribution for $\left(\tilde{X}_{res}^{*},\tilde{\epsilon}_{res},\tilde{U}_{1},\tilde{U}_{2}\right)$ is a valid normal distribution.
Second, we highlight that the candidate latent variables satisfy Equations (C.20)-(C.24) by construction.
Lastly, we must show that the data restrictions (Equations (C.25)-(C.33)) are valid. Since we have shown that candidate latent variables satisfy Equations (C.20)-(C.33), we can follow the same steps used to prove Lemma C.1 to show that
$\mathbb{E}\left[\tilde{Y}_{res}\right]$ $=0$ (C.48)
$\mathbb{E}\left[\tilde{Z}_{1,res}\right]$ $=\mathbb{E}\left[\tilde{U}_{1}\right]$ (C.49)
$\mathbb{E}\left[\tilde{Z}_{2,res}\right]$ $=\mathbb{E}\left[\tilde{U}_{2}\right]$ (C.50)
$Var\left(\tilde{Y}_{res}\right)$ $=\tilde{\beta}^{2}+Var\left(\tilde{\epsilon}_{res}\right)$ (C.51)
$Var\left(\tilde{Z}_{1,res}\right)$ $=1+Var\left(\tilde{U}_{1}\right)+2\cdot Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{1}\right)$ (C.52)
$Var\left(\tilde{Z}_{2,res}\right)$ $=1+Var\left(\tilde{U}_{2}\right)+2\cdot Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{2}\right)$ (C.53)
$Cov\left(\tilde{Z}_{1,res},\tilde{Y}_{res}\right)$ $=\tilde{\beta}+\tilde{\beta}\cdot Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{1}\right)$ (C.54)
$Cov\left(\tilde{Z}_{2,res},\tilde{Y}_{res}\right)$ $=\tilde{\beta}+\tilde{\beta}\cdot Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{2}\right)$ (C.55)
$Cov\left(\tilde{Z}_{1,res},\tilde{Z}_{2,res}\right)$ $=1+Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{1}\right)+Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{2}\right)+Cov\left(\tilde{U}_{1},\tilde{U}_{2}\right).$ (C.56)
Based on Equation (C.48), we have that
$\mathbb{E}\left[\tilde{Y}_{res}\right]$ $=0=\mathbb{E}\left[Y_{res}\right]$
according to Equation (C.4), implying that Equation (C.25) is valid.
Based on Equation (C.49), we have that
$\mathbb{E}\left[\tilde{Z}_{1,res}\right]$ $=\mathbb{E}\left[\tilde{U}_{1}\right]=\mathbb{E}\left[Z_{1,res}\right]$
according to Equation (C.36), implying that Equation (C.26) is valid.
Based on Equation (C.50), we have that
$\mathbb{E}\left[\tilde{Z}_{2,res}\right]$ $=\mathbb{E}\left[\tilde{U}_{2}\right]=\mathbb{E}\left[Z_{2,res}\right]$
according to Equation (C.37), implying that Equation (C.27) is valid.
Based on Equation (C.51), we have that
$Var\left(\tilde{Y}_{res}\right)$ $=\tilde{\beta}^{2}+Var\left(\tilde{\epsilon}_{res}\right)$
$=\tilde{\beta}^{2}+\left[Var\left(Y_{res}\right)-\tilde{\beta}^{2}\right]$
by Equation (C.39)
$=Var\left(Y_{res}\right),$
implying that Equation (C.28) is valid.
Based on Equation (C.52), we have that
$Var\left(\tilde{Z}_{1,res}\right)$
$\hskip 20.0pt=1+Var\left(\tilde{U}_{1}\right)+2\cdot Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{1}\right)$
$\hskip 20.0pt=1+\left[1+Var\left(Z_{1,res}\right)-2\cdot\dfrac{Cov\left(Z_{1,res},Y_{res}\right)}{\tilde{\beta}}\right]+2\cdot\left[\dfrac{Cov\left(Z_{1,res},Y_{res}\right)}{\tilde{\beta}}-1\right]$
by Equations (C.40) and (C.45)
$\hskip 20.0pt=Var\left(Z_{1,res}\right),$
implying that Equation (C.29) is valid.
Based on Equation (C.53), we have that
$Var\left(\tilde{Z}_{2,res}\right)$
$\hskip 20.0pt=1+Var\left(\tilde{U}_{2}\right)+2\cdot Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{2}\right)$
$\hskip 20.0pt=1+\left[1+Var\left(Z_{2,res}\right)-2\cdot\dfrac{Cov\left(Z_{2,res},Y_{res}\right)}{\tilde{\beta}}\right]+2\cdot\left[\dfrac{Cov\left(Z_{2,res},Y_{res}\right)}{\tilde{\beta}}-1\right]$
by Equations (C.41) and (C.46)
$\hskip 20.0pt=Var\left(Z_{2,res}\right),$
implying that Equation (C.30) is valid.
Based on Equation (C.54), we have that
$Cov\left(\tilde{Z}_{1,res},\tilde{Y}_{res}\right)$
$\hskip 20.0pt=\tilde{\beta}+\tilde{\beta}\cdot Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{1}\right)$
$\hskip 20.0pt=\tilde{\beta}+\tilde{\beta}\cdot\left[\dfrac{Cov\left(Z_{1,res},Y_{res}\right)}{\tilde{\beta}}-1\right]$
by Equation (C.45)
$\hskip 20.0pt=Cov\left(Z_{1,res},Y_{res}\right),$
implying that Equation (C.31) is valid.
Based on Equation (C.55), we have that
$Cov\left(\tilde{Z}_{2,res},\tilde{Y}_{res}\right)$
$\hskip 20.0pt=\tilde{\beta}+\tilde{\beta}\cdot Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{2}\right)$
$\hskip 20.0pt=\tilde{\beta}+\tilde{\beta}\cdot\left[\dfrac{Cov\left(Z_{2,res},Y_{res}\right)}{\tilde{\beta}}-1\right]$
by Equation (C.46)
$\hskip 20.0pt=Cov\left(Z_{2,res},Y_{res}\right),$
implying that Equation (C.32) is valid.
Based on Equation (C.56), we have that
$Cov\left(\tilde{Z}_{1,res},\tilde{Z}_{2,res}\right)$
$\hskip 20.0pt=1+Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{1}\right)+Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{2}\right)+Cov\left(\tilde{U}_{1},\tilde{U}_{2}\right)$
$\hskip 20.0pt=1+\left[\dfrac{Cov\left(Z_{1,res},Y_{res}\right)}{\tilde{\beta}}-1\right]+\left[\dfrac{Cov\left(Z_{2,res},Y_{res}\right)}{\tilde{\beta}}-1\right]$
[…]
by Equations (C.45)-(C.47)
$=Cov\left(Z_{1,res},Z_{2,res}\right),$
implying that Equation (C.33) is valid.
This concludes the proof for the case $\tilde{\beta}>0$ because we have constructed a distribution for the candidate latent variables that satisfies the model restrictions and the data restrictions.
Part 2: $\tilde{\beta}=0$
First, we note that $\tilde{\beta}=0$ implies that $LB_{C}=0$ because $LB_{C}\geq 0$ by Assumption C.5. Consequently, we have that
$Cov\left(Z_{1,res},Y_{res}\right)=0$ (C.57)
and
$Cov\left(Z_{2,res},Y_{res}\right)=0.$ (C.58)
Now, we need to create a distribution for the candidate latent variables $\left(\tilde{X}_{res}^{*},\tilde{\epsilon}_{res},\tilde{U}_{1},\tilde{U}_{2}\right)$. The candidate observable variables $\left(\tilde{Y}_{res},\tilde{Z}_{1,res},\tilde{Z}_{1,res}\right)$ are created based on Equations (C.17)-(C.19).
We impose
$\left(\tilde{X}_{res}^{*},\tilde{\epsilon}_{res},\tilde{U}_{1},\tilde{U}_{2}\right)\sim N\left(\boldsymbol{\mu},\boldsymbol{\Omega}\right),$
i.e., $\left(\tilde{X}_{res}^{*},\tilde{\epsilon}_{res},\tilde{U}_{1},\tilde{U}_{2}\right)$ are jointly normally distributed with mean $\boldsymbol{\mu}$ and covariance matrix $\boldsymbol{\Omega}$.
The vector $\boldsymbol{\mu}$ is given by
$\mathbb{E}\left[\tilde{X}_{res}^{*}\right]$ $\coloneqq 0$ (C.59)
$\mathbb{E}\left[\tilde{\epsilon}_{res}\right]$ $\coloneqq 0$ (C.60)
$\mathbb{E}\left[\tilde{U}_{1}\right]$ $\coloneqq\mathbb{E}\left[Z_{1,res}\right]$ (C.61)
$\mathbb{E}\left[\tilde{U}_{1}\right]$ $\coloneqq\mathbb{E}\left[Z_{2,res}\right]$ (C.62)
and the covariance matrix $\boldsymbol{\Omega}$ is given by
$Var\left(\tilde{X}_{res}^{*}\right)$ $\coloneqq 1$ (C.63)
$Var\left(\tilde{\epsilon}_{res}\right)$ $\coloneqq Var\left(Y_{res}\right)$ (C.64)
$Var\left(\tilde{U}_{1}\right)$ $\coloneqq 1+Var\left(Z_{1,res}\right)$ (C.65)
$Var\left(\tilde{U}_{2}\right)$ $\coloneqq 1+Var\left(Z_{2,res}\right)$ (C.66)
$Cov\left(\tilde{X}_{res}^{*},\tilde{\epsilon}_{res}\right)$ $\coloneqq 0$ (C.67)
$Cov\left(\tilde{U}_{1},\tilde{\epsilon}_{res}\right)$ $\coloneqq 0$ (C.68)
$Cov\left(\tilde{U}_{2},\tilde{\epsilon}_{res}\right)$ $\coloneqq 0$ (C.69)
$Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{1}\right)$ $\coloneqq-1$ (C.70)
$Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{2}\right)$ $\coloneqq-1$ (C.71)
$Cov\left(\tilde{U}_{1},\tilde{U}_{2}\right)$ $\coloneqq Cov\left(Z_{1,res},Z_{2,res}\right)+1.$ (C.72)
Observe that the above distribution is a valid normal distribution because $Var\left(\tilde{\epsilon}_{res}\right)\geq 0$, $Var\left(\tilde{U}_{1}\right)\geq 0$ and $Var\left(\tilde{U}_{2}\right)\geq 0$ by construction.
Now, notice that the candidate latent variables satisfy Equations (C.20)-(C.24) by construction.
Lastly, we must show that the data restrictions (Equations (C.25)-(C.33)) are valid. Since we have shown that candidate latent variables satisfy Equations (C.20)-(C.24), we can follow the same steps used to prove Lemma C.1 to show that
$\mathbb{E}\left[\tilde{Y}_{res}\right]$ $=0$ (C.73)
$\mathbb{E}\left[\tilde{Z}_{1,res}\right]$ $=\mathbb{E}\left[\tilde{U}_{1}\right]$ (C.74)
$\mathbb{E}\left[\tilde{Z}_{2,res}\right]$ $=\mathbb{E}\left[\tilde{U}_{2}\right]$ (C.75)
$Var\left(\tilde{Y}_{res}\right)$ $=\tilde{\beta}^{2}+Var\left(\tilde{\epsilon}_{res}\right)$ (C.76)
$Var\left(\tilde{Z}_{1,res}\right)$ $=1+Var\left(\tilde{U}_{1}\right)+2\cdot Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{1}\right)$ (C.77)
$Var\left(\tilde{Z}_{2,res}\right)$ $=1+Var\left(\tilde{U}_{2}\right)+2\cdot Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{2}\right)$ (C.78)
$Cov\left(\tilde{Z}_{1,res},\tilde{Y}_{res}\right)$ $=\tilde{\beta}+\tilde{\beta}\cdot Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{1}\right)$ (C.79)
$Cov\left(\tilde{Z}_{2,res},\tilde{Y}_{res}\right)$ $=\tilde{\beta}+\tilde{\beta}\cdot Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{2}\right)$ (C.80)
$Cov\left(\tilde{Z}_{1,res},\tilde{Z}_{2,res}\right)$ $=1+Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{1}\right)+Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{2}\right)+Cov\left(\tilde{U}_{1},\tilde{U}_{2}\right).$ (C.81)
Based on Equation (C.73), we have that
$\mathbb{E}\left[\tilde{Y}_{res}\right]$ $=0=\mathbb{E}\left[Y_{res}\right]$
according to Equation (C.4), implying that Equation (C.25) is valid.
Based on Equation (C.74), we have that
$\mathbb{E}\left[\tilde{Z}_{1,res}\right]$ $=\mathbb{E}\left[\tilde{U}_{1}\right]=\mathbb{E}\left[Z_{1,res}\right]$
according to Equation (C.61), implying that Equation (C.26) is valid.
Based on Equation (C.75), we have that
$\mathbb{E}\left[\tilde{Z}_{2,res}\right]$ $=\mathbb{E}\left[\tilde{U}_{2}\right]=\mathbb{E}\left[Z_{2,res}\right]$
according to Equation (C.62), implying that Equation (C.27) is valid.
Based on Equation (C.76), we have that
$Var\left(\tilde{Y}_{res}\right)$ $=Var\left(\tilde{\epsilon}_{res}\right)=Var\left(Y_{res}\right)$
by Equation (C.64), implying that Equation (C.28) is valid.
Based on Equation (C.77), we have that
$Var\left(\tilde{Z}_{1,res}\right)$
$\hskip 20.0pt=1+Var\left(\tilde{U}_{1}\right)+2\cdot Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{1}\right)$
$\hskip 20.0pt=1+\left[1+Var\left(Z_{1,res}\right)\right]+2\cdot\left[-1\right]$
by Equations (C.65) and (C.70)
$\hskip 20.0pt=Var\left(Z_{1,res}\right),$
implying that Equation (C.29) is valid.
Based on Equation (C.78), we have that
$Var\left(\tilde{Z}_{2,res}\right)$
$\hskip 20.0pt=1+Var\left(\tilde{U}_{2}\right)+2\cdot Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{2}\right)$
$\hskip 20.0pt=1+\left[1+Var\left(Z_{2,res}\right)\right]+2\cdot\left[-1\right]$
by Equations (C.66) and (C.71)
$\hskip 20.0pt=Var\left(Z_{2,res}\right),$
implying that Equation (C.30) is valid.
Based on Equation (C.79), we have that
$Cov\left(\tilde{Z}_{1,res},\tilde{Y}_{res}\right)$ $=0=Cov\left(Z_{1,res},Y_{res}\right)$
by Equation (C.57), implying that Equation (C.31) is valid.
Based on Equation (C.80), we have that
$Cov\left(\tilde{Z}_{2,res},\tilde{Y}_{res}\right)$ $=0=Cov\left(Z_{2,res},Y_{res}\right)$
by Equation (C.58), implying that Equation (C.32) is valid.
Based on Equation (C.56), we have that
$Cov\left(\tilde{Z}_{1,res},\tilde{Z}_{2,res}\right)$
$\hskip 20.0pt=1+Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{1}\right)+Cov\left(\tilde{X}_{res}^{*},\tilde{U}_{2}\right)+Cov\left(\tilde{U}_{1},\tilde{U}_{2}\right)$
$\hskip 20.0pt=1+\left[-1\right]+\left[-1\right]+\left[1+Cov\left(Z_{1,res},Z_{2,res}\right)\right]$
by Equations (C.70)-(C.72)
$=Cov\left(Z_{1,res},Z_{2,res}\right),$
implying that Equation (C.33) is valid.
This concludes the proof for the case $\tilde{\beta}=0$ because we have constructed a distribution for the candidate latent variables that satisfies the model restrictions and the data restrictions.
Combining Parts 1 and 2 of Appendix C.1.4, we can conclude that, if $UB\geq LB_{C}$, then the interval $\left[LB_{C},UB\right]$ exhausts all the information contained in the first two moments of the data distribution.
Appendix D No Prior Knowledge of the Target Parameter’s Sign
In this appendix, we propose a partial identification strategy that requires no prior knowledge about the sign of the true target parameter $\beta$. To do so, we keep Assumptions 1, 2, 4 and 5, and modify Assumptions 3 and 6. For convenience, we list all the necessary assumptions here.
Assumption D.1 (Exogeneity)
After conditioning on the covariate vector $W$, the true unobserved measure $X^{*}$ is exogenous, i.e., $Cov\left(X_{res}^{*},\epsilon_{res}\right)=0$.
Assumption D.2 (Scale Normalization)
The scale of the true unobserved measure is normalized to have unit variance, i.e., $Var\left(X_{res}^{*}\right)=1$.
Assumption D.3 (Reduced-Form Coefficients)
The reduced-form coefficients identify the sign of the true target parameter, i.e., $sign\left\{Cov\left(Z_{1},Y_{res}\right)\right\}=sign\left\{Cov\left(Z_{2},Y_{res}\right)\right\}=sign\left\{\beta\right\}$. [^17] The signsign function equals one if its argument is a strictly positive number, zero if its argument is zero, and minus one if its argument is a strictly negative number.
Assumption D.4 (Exogeneity of the Measurement Errors)
The measurement error variables are exogenous, i.e., $Cov\left(U_{1},\epsilon_{res}\right)=0$ and $Cov\left(U_{2},\epsilon_{res}\right)=0$.
Assumption D.5 (Nonclassical Measurement Error)
The measurement error terms are positively correlated with each other, i.e., $Cov\left(U_{1},U_{2}\right)\geq 0$.
Assumption D.6 (Well-behaved Data Distributions)
The variance of the outcome variable is strictly positive, and the covariance of the two proxy variables is strictly greater than the negative of the variable of interest’s variance, i.e., $Var\left(Y_{res}\right)>0$ and $Cov\left(Z_{1,res},Z_{2,res}\right)>-1.$
We now state our main result: partial identification of $\beta$ (Equation (2)) using two proxies (Equation (3)) with nonclassical measurement error and no prior knowledge of target parameter’s sign.
Proposition D.1
Under Assumptions D.1-D.6, we have that:
$\beta\in\left\{\begin{array}[]{cl}\left\{\emptyset\right\}&\text{ if }UB<LB\\ \left[LB,UB\right]&\text{ if }UB\geq LB\end{array}\right$. (D.1)
if $LB>0$,
$\beta=0\text{ if }LB=0$
and
$\beta\in\left\{\begin{array}[]{cl}\left\{\emptyset\right\}&\text{ if }LB<-UB\\ \left[-UB,LB\right]&\text{ if }LB\geq-UB\end{array}\right$. (D.2)
if $LB<0$.
Moreover, these bounds exploit all the information contained in the first two moments of the observable variables $\left(Y_{res},Z_{1,res},Z_{2,res}\right)$.
Proof.
See Appendix D.1. ∎
The main drawback of Proposition D.1 in comparison to Proposition 1 is that this appendix’s result requires testing the sign of $LB$ before estimating the bounds around $\beta$. This pre-testing step causes some challenges in controlling size when conducting inference.
Proof of Proposition D.1
When $LB>0$, Assumption D.3 implies that $\beta>0$. We can then follow the steps described in Appendix A.1 to derive Equation (D.1) and to show that these bounds are sharp.
When $LB=0$, Assumption D.3 implies that $\beta=0$ and we have point-identification.
When $LB<0$, Assumption D.3 implies that $\beta<0$. To derive Equation D.2, we start by rewriting Equation (2) as
$\check{Y}_{res}=\check{\beta}\cdot X_{res}^{*}+\check{\epsilon}_{res},$
where $\check{Y}_{res}\coloneqq-Y_{res}$, $\check{\beta}\coloneqq-\beta$ and $\check{\epsilon}_{res}\coloneqq-\epsilon_{res}$. Since $\check{\beta}>0$, we can follow the steps described in Appendix A.1 to derive Equation (D.2) and to show that these bounds are sharp.
Appendix E Estimation and Inference: Detailed Explanation
In this appendix, we propose a simple parametric estimator for the bounds described in Proposition 1 (Appendix E.1) and a method to compute $\alpha$-confidence sets that contain the identified set with a probability at least as larger as the chosen confidence level $\alpha\in\left(0,1\right)$ (Appendix E.2).
Throughout this appendix, we assume that we observe an independently and identically distributed sample $\left\{Y_{i},W_{1,i},\ldots,W_{J,i},Z_{1,i},Z_{2,i}\right\}_{i=1}^{N}$, where $N\in\mathbb{N}$ is the sample size.
Estimation
In this section, we propose a simple parametric estimator for the bounds described in Proposition 1. Algorithm 1 describes how to implement this estimator in nine steps.
Algorithm 1 (Parametric Estimator for the Identified Set)
a
  1. 1.
    Run the OLS regression
    $Y_{i}=a_{0}^{Y}+\sum_{j=1}^{J}a_{j}^{Y}\cdot W_{j,i}+\epsilon_{i}^{Y},$
    where $\widehat{a}^{Y}\coloneqq\left(\widehat{a}_{0}^{Y},\widehat{a}_{1}^{Y},\cdots,\widehat{a}_{J}^{Y}\right)$ are the OLS estimators, and save the residuals
    $\widehat{Y}_{res,i}\coloneqq Y_{i}-\widehat{a}_{0}^{Y}-\sum_{j=1}^{J}\widehat{a}_{j}^{Y}\cdot W_{j,i}.$
  2. 2.
    Run the OLS regression
    $Z_{1,i}=a_{0}^{1}+\sum_{j=1}^{J}a_{j}^{1}\cdot W_{j,i}+\epsilon_{i}^{1},$
    where $\widehat{a}^{1}\coloneqq\left(\widehat{a}_{0}^{1},\widehat{a}_{1}^{1},\cdots,\widehat{a}_{J}^{1}\right)$ are the OLS estimators, and save the residuals
    $\widehat{Z}_{1,res,i}\coloneqq Z_{1,i}-\widehat{a}_{0}^{1}-\sum_{j=1}^{J}\widehat{a}_{j}^{1}\cdot W_{j,i}.$
  3. 3.
    Run the OLS regression
    $Z_{2,i}=a_{0}^{2}+\sum_{j=1}^{J}a_{j}^{2}\cdot W_{j,i}+\epsilon_{i}^{2},$
    where $\widehat{a}^{2}\coloneqq\left(\widehat{a}_{0}^{2},\widehat{a}_{2}^{2},\cdots,\widehat{a}_{J}^{2}\right)$ are the OLS estimators, and save the residuals
    $\widehat{Z}_{2,res,i}\coloneqq Z_{2,i}-\widehat{a}_{0}^{2}-\sum_{j=1}^{J}\widehat{a}_{j}^{2}\cdot W_{j,i}.$
  4. 4.
    Estimate
    $\widehat{l}_{1}\coloneqq\dfrac{Cov\left(\widehat{Z}_{1,res,i},\widehat{Y}_{res,i}\right)+Cov\left(\widehat{Z}_{2,res,i},\widehat{Y}_{res,i}\right)}{1+Cov\left(\widehat{Z}_{1,res,i},\widehat{Z}_{2,res,i}\right)}.$ (E.1)
  5. 5.
    Estimate
    $\widehat{l}_{2}\coloneqq 2\cdot\dfrac{Cov\left(\widehat{Z}_{1,res,i},\widehat{Y}_{res,i}\right)}{1+Var\left(\widehat{Z}_{1,res,i}\right)}.$ (E.2)
  6. 6.
    Estimate
    $\widehat{l}_{3}\coloneqq 2\cdot\dfrac{Cov\left(\widehat{Z}_{2,res,i},\widehat{Y}_{res,i}\right)}{1+Var\left(\widehat{Z}_{2,res,i}\right)}.$ (E.3)
  7. 7.
    Estimate
    $\widehat{LB}\coloneqq\max\left\{\widehat{l}_{1},\widehat{l}_{2},\widehat{l}_{3}\right\}.$ (E.4)
  8. 8.
    Estimate
    $\widehat{UB}\coloneqq\sqrt{Var\left(\widehat{Y}_{res}\right)}.$ (E.5)
  9. 9.
    If $\widehat{UB}<\widehat{LB}$, then the estimator for the identified set in Proposition 1 is given by $\emptyset$. If $\widehat{UB}\geq\widehat{LB}$, then the estimator for the identified set in Proposition 1 is given by $\left[\widehat{LB},\widehat{UB}\right$].
Inference
In this section, we address the sample uncertainty behind the estimation of our bounds. In summary, we propose one method to construct an $\alpha$-confidence interval around the identified set, $\left[LB,UB\right]$. Specifically, this interval contains the true identified set with known probability, where the probability is computed based on the thought exercise of an infinitely large number of samples.
To construct our $\alpha$-confidence interval around the identified set, we combine the confidence bounds proposed by Chernozhukov et al. (2013) with a Bonferroni-style correction. Since the upper bound does not rely on intersection bounds, we use a simple bootstrap procedure to compute a $\frac{\left(1+\alpha\right)}{2}$-confidence bound above it. Since the lower bound relies on intersection bounds, we use the method proposed by Chernozhukov et al. (2013) to compute a $\frac{\left(1+\alpha\right)}{2}$-confidence bound below it. By taking the intersection of these two confidence bounds, we construct an $\alpha$-confidence interval around the identified set. Similar procedures were used by Flores and Flores-Lagunes (2013) and Possebom and Riva (2024).
Algorithm 2 describes how to construct our confidence intervals.
Algorithm 2 ($\alpha$-Confidence Interval around the Identified Set)
a
  1. 1.
    Upper $\frac{\left(1+\alpha\right)}{2}$-Confidence Bound
    1. (a)
      Using the empirical bootstrap, collect $B\in\mathbb{N}$ bootstrapped estimates of the upper bound, $\left\{\widehat{UB}_{b}\right\}_{b=1}^{B}$.
    2. (b)
      Obtain the $\frac{\left(1+\alpha\right)}{2}$-quantile of $\left\{\left|\widehat{UB}_{b}-\widehat{UB}\right|\right\}_{b=1}^{B}$ and denote it by $c_{UB}\left(\frac{1+\alpha}{2}\right)$.
    3. (c)
      Construct the Upper $\frac{\left(1+\alpha\right)}{2}$-Confidence Bound as $\widehat{C}_{UB}\left(\frac{1+\alpha}{2}\right)\coloneqq\widehat{UB}+c_{UB}\left(\frac{1+\alpha}{2}\right)$.
  2. 2.
    Lower $\frac{\left(1+\alpha\right)}{2}$-Confidence Bound
    1. (a)
      Define the vector of bounding functions,
      $l\coloneqq\left(\dfrac{Cov\left(Z_{1,res},Y_{res}\right)+Cov\left(Z_{2,res},Y_{res}\right)}{1+Cov\left(Z_{1,res},Z_{2,res}\right)},2\cdot\dfrac{Cov\left(Z_{1,res},Y_{res}\right)}{1+Var\left(Z_{1,res}\right)},2\cdot\dfrac{Cov\left(Z_{2,res},Y_{res}\right)}{1+Var\left(Z_{2,res}\right)}\right)^{\prime},$
      and its estimated version, $\widehat{l}\coloneqq\left(\widehat{l}_{1},\widehat{l}_{2},\widehat{l}_{3}\right)^{\prime}$.
    2. (b)
      Using the empirical bootstrap, collect $B\in\mathbb{N}$ bootstrapped estimates of the bounding function, $\left\{\widehat{l}_{b}^{*}\right\}_{b=1}^{B}$.
    3. (c)
      Estimate the variance-covariance matrix of $\sqrt{N}\cdot\left(\widehat{l}-l\right)$ using the variance-covariance matrix of $\left\{\sqrt{N}\cdot\widehat{l}_{b}^{*}\right\}_{b=1}^{B}$. Denote the estimated variance-covariance matrix by $\widehat{\Omega}$.
    4. (d)
      Let $\widehat{g}_{k}^{\prime}$ denote the $k$-th row of $\widehat{\Omega}^{\nicefrac{{1}}{{2}}}$ for $k\in\left\{1,2,3\right\}$.
    5. (e)
[…]
    6. (f)
      Simulate $R\in\mathbb{N}$ draws from the multivariate standard normal of dimension 3 and denote them by $Z_{1},\ldots,Z_{R}$ as row vectors.
    7. (g)
[…]
    8. (h)
      Let $Q_{p}\left(A\right)$ denote the $p$-th quantile of a random variable A and, following Chernozhukov et al. (2013), define $c_{N}\coloneqq 1-\frac{0.1}{\ln{N}}$.
    9. (i)
      Compute $\kappa\coloneqq Q_{c_{N}}\left(\max\left\{Z_{r}^{*}\left(1\right),Z_{r}^{*}\left(2\right),Z_{r}^{*}\left(3\right)\right\};r=1,\ldots,R\right)$. In plain English, we calculate the maximum of $Z_{r}^{*}\left(1\right),Z_{r}^{*}\left(2\right)$ and $Z_{r}^{*}\left(3\right)$ for each replication $r$, and take the $c_{N}$-th quantile of those $R$ values.
    10. (j)
[…]
    11. (k)
      Compute $\widehat{\kappa}\coloneqq Q_{\frac{\left(1+\alpha\right)}{2}}\left(\max\limits_{k\in\widehat{K}}\left\{Z_{r}^{*}\left(k\right)\right\};r=1,\ldots,R\right)$.
    12. (l)
      Construct the Lower $\frac{\left(1+\alpha\right)}{2}$-Confidence Bound as
      $\widehat{C}_{LB}\left(\frac{1+\alpha}{2}\right)\coloneqq\max\limits_{k\in\left\{1,2,3\right\}}\left\{\widehat{l}_{k}-\widehat{\kappa}\cdot\widehat{s}_{k}\right\}.$
  3. 3.
    $\alpha$-Confidence Interval
    1. (a)
      Construct a $\alpha$-Confidence Interval using $\left[\widehat{C}_{LB},\widehat{C}_{UB}\right]$. This interval contains the true identified set $\left[LB,UB\right]$ with probability $\alpha\in\left(0.5,1\right)$.
Appendix F Applications Using Other Datasets
In this appendix, we discuss other empirical contexts where our partial identification strategy may be useful. To illustrate the broad applicability of our methodology, we will briefly analyze six empirical examples where the variable of interest is measured with error, and there might exist two possible proxies.
First, many recent studies (Olken, 2009; Adena et al., 2015) focus on the impact of media access on economic behavior. To do so, they frequently measure media access by combining an irregular terrain model and the location of TV or radio transmitters. Since the most commonly used irregular terrain model is in its third version (Hufford, 1995), this setting is very similar to the FAO-GAEZ productivity measures discussed in Sections 2.1 and 5. For example, we can use our proposed tool to analyze a reduced-form version of Panel B in Table 11 by Olken (2009) and reanalyze Equation (1) by Adena et al. (2015).
Second, many articles (Post and Hanewald, 2013; Gan et al., 2015; Bissonnette et al., 2017) discuss the impact of subjective beliefs about life expectancy on economic behavior (e.g., savings, investment, bequests). To measure this type of belief, many surveys ask individuals about their subjective probability of attaining various ages. However, individual responses frequently present measurement issues, such as rounding and focal answers. To address these problems, many authors adopt different imputation methods that approximate the true subjective probability of attaining a certain age. By treating the results of each imputation model as a different proxy, our proposed method may be used in this empirical setting. For example, we can reanalyze Equation (4-3) by Post and Hanewald (2013).
Third, multiple studies (Hsieh and Klenow, 2009, 2010) are interested in the impact of total factor productivity on economic outcomes, such as plant exit and income levels. To measure total factor productivity, different authors adopt different models (Olley and Pakes, 1996; Hsieh and Klenow, 2009; Ackerberg et al., 2015). By treating the measures produced by each model as a different proxy for the true latent total factor productivity, our partial identification strategy may be used in this empirical context. For example, we can reanalyze Table VIII by Hsieh and Klenow (2009).
Fourth, economic historians (Sala-i-Martin, 1996; Hartley, 2021) frequently focus on understanding GDP growth and fluctuations in the long run. To do so, they need measures of income levels over long periods of time, including moments when national accounts were not collected in a systematic way. To access this type of variable, most authors use the Maddison Project Data (Bolt and van Zanden, 2025), which is in its fifth version. Each version varies either because new countries were included in the sample or, more importantly to us, new information was collected about existing countries and their historical income data were updated. Since different versions of these variables may have different measurement error terms when compared to the true income level of each country, this setting is very similar to the FAO-GAEZ productivity measures discussed in Sections 2.1 and 5. For example, we can use our proposed tool to reanalyze Equations (1), (3) and (5) by Sala-i-Martin (1996) and Equations (1) and (2) by Hartley (2021).
Fifth, Rodrik and Wacziarg (2005, Tables 1 and 2) analyze the impact of democratic transitions on economic outcomes. To measure their treatment variable, they use the Polity IV index. This expert-coded measure of democracy levels is in its fourth version and is based on experts’ interpretation of political regimes in each country and each year. Since different experts might interpret the same political regime differently, there is measurement error in each version of this index. Although this setting is similar to the FAO-GAEZ productivity measures (Sections 2.1 and 5), we believe that Assumption 5 may not be appropriate in this context. Since political experts might overreact to reanalyses of historical phenomena, it may be the case that the measurement error terms in different versions of the Polity Index are negatively correlated. In this case, the bounds proposed in Appendix C represent a more appropriate approach than the bounds in the main text.
Sixth, multiple articles (Alfaro et al., 2008; Rajkumar and Swaroop, 2008; Aksoy et al., 2025) are interested in the impact of a country’s governance quality on economic outcomes, such as health quality, education quality, political trust and international capital flows. To measure governance quality, Alfaro et al. (2008) and Rajkumar and Swaroop (2008) use a variety of measures from the Political Risk Services Group, while Aksoy et al. (2025) uses the V-DEM index and one measure from the Worldwide Governance Indicators. By treating these different measures as different proxies for the true unobservable governance quality, our partial identification strategy may be applied in this empirical context. For example, we can reanalyze Tables 5, 9 and 10 (Panel C) by Alfaro et al. (2008), Equations (1) and (2) by Rajkumar and Swaroop (2008), and Equation (2) by Aksoy et al. (2025).
Appendix G A Discussion on Partial Identification
In this appendix, we intuitively explain the concept of partial identification and why it is a useful tool for applied researchers. For deeper reviews about partial identification, the interested reader may check the work by Tamer (2010), Manski (2011), Canay and Shaikh (2017), and Lewbel (2018). To understand the importance of partial identification, we first generically define a data-generating process, a target parameter and point identification. Then, we intuitively define model uncertainty and partial identification, explaining how the latter addresses the former. Lastly, we discuss sample uncertainty and confidence intervals in a context of partially identified target parameters.
A data-generating process is a statistical model. It defines the distribution of latent variables and the functions that connect latent variables to observable variables. These functions may be parametric or non-parametric. For example, in our context, the data-generating process includes the coefficients of a linear regression model, the true unobservable productivity, the economic shock in the outcome equation, and the measurement error terms in the two versions of the FAO-GAEZ proxies.
Importantly, the data-generating process includes the model assumptions through restrictions in the distribution of the latent variables. In our context, it imposes that the economic shock is uncorrelated with the true unobservable productivity and with the measurement error terms. It also imposes that the two measurement error terms are positively correlated with each other.
A target parameter is a function of the data-generating process. In our context, it is the coefficient of a linear regression and captures the effect of increasing agricultural productivity by one standard deviation.
A point-identification result shows that a target parameter is equal to a function of the distribution of observable random variables. [^18] To fix ideas, imagine a randomized control trial with perfect compliance. The data-generating process defines the potential outcomes, the treatment assignment variable and the usual definition of the observed outcome. The target parameter is the expected value of the difference between the treated and untreated potential outcomes. In particular, this target parameter is identified by the difference between the expected observed outcome of treated individuals and the expected observed outcome of control individuals. Importantly, all objects here are defined at the population level. In our context, we could have derived a point-identification result by imposing that there is no measurement error. In this case, we would observe the true productivity and could point-identify its effect by regressing the outcome variable on our measure of agricultural productivity.
The point-identification approach usually requires strong restrictions on the distribution of the latent variables (Manski, 2011). However, researchers may be unsure about the true model of the data-generating process. Two equally reasonable researchers may disagree about the plausibility of the assumptions included in the data-generating process. They may even believe in different sets of assumptions (Tamer, 2010).
Partial identification tools address this type of model uncertainty. A scalar target parameter is partially identified if there is an interval that contains the target parameter and whose boundary points are functions of the distribution of observable random variables. A researcher may adopt a weaker set of assumptions, which is not sufficient for point-identification but is sufficient for partial identification. The length of this identified interval is an intuitive measure of model uncertainty. If there is less model uncertainty, the interval is shorter and may even collapse to a point.
Moreover, a different researcher may adopt an even weaker set of assumptions, which still partially identifies the target parameter. However, due to imposing even weaker assumptions, the identified set is wider. This extra length is intuitively connected with the additional model uncertainty, which is represented by weaker assumptions.
Our approach allows for both types of researchers. The first researcher (Section 3) is not willing to assume that measurement error does not exist but is willing to impose restrictions on the sign of the true effect and on the sign of the correlation between the measurement error terms. The second researcher (Appendices C and D) is not willing to make one of these assumptions and finds a wider identified set.
Lastly, bounds in partial identification results are defined at the population level. Consequently, they do not take into account sample uncertainty, i.e., they do not consider that a finite sample is a random draw from a superpopulation. Similarly to the point-identification approach, the partial identification approach takes sample uncertainty into account when estimating the bounds of the identified set. It does so by estimating confidence intervals.
There are two types of confidence intervals in the partial identification approach. The first type of confidence interval contains the true target parameter with a known probability, where the probability is computed based on the thought experiment of an infinitely large number of samples. The second type of confidence interval contains the entire identified set with a known probability. The trade-off between these types of confidence intervals is that the second confidence interval is usually wider but easier to estimate. In our context, Section 4 explains how to construct the second type of confidence interval.
Appendix H Assumptions on Model Primitives
In this appendix, we describe five assumptions on the primitives of our model that imply Assumptions 1-5 in Section 3.1. The primitives of our model are the random variables $Y$, $W_{1}$, …, $W_{J}$, $X^{*}$, $\epsilon$, $Z_{1}$ and $Z_{2}$, and the following outcome equation:
$Y=\alpha_{0}+\sum_{j=1}^{J}\alpha_{j}\cdot W_{j}+\beta\cdot X^{*}+\epsilon,$ (H.1)
Using the random variables in our model primitives, we define five additional random variables:
$X_{res}^{*}\coloneqq X^{*}-W\cdot\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot X^{*}\right],$ (H.2)
$Z_{k,res}\coloneqq Z_{k}-W\cdot\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot Z_{k}\right]$ (H.3)
and
$U_{k}\coloneqq Z_{k,res}-X_{res}^{*}$ (H.4)
for any $k\in\left\{1,2\right\}$, where $W=\left[1,W_{1},\ldots,W_{J}\right]$ is the row vector of covariates including the constant term.
Our first assumption on the primitives of our model imposes that the economic shock is exogenous.
Assumption H.1
$Cov\left(A,\epsilon\right)=0$ for any $A\in\left\{W_{1},\ldots,W_{J},X^{*}\right\}$.
This assumption implies that
$\epsilon_{res}$ $\coloneqq\epsilon-W\cdot\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot\epsilon\right]$
$=\epsilon$
and that
$Cov\left(X_{res}^{*},\epsilon_{res}\right)$ $=Cov\left(X^{*}-W\cdot\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot X^{*}\right],\epsilon\right)$
$=Cov\left(X^{*},\epsilon\right)-Cov\left(W,\epsilon\right)\cdot\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot X^{*}\right]$
$=0.$
Hence, Assumption H.1 implies Assumption 1.
Our second assumption on the primitives of our model is Assumption 2 directly because we see it as modifying the target parameter, as explained in the main text.
Our third assumption on the primitives of our model is Assumption 3 directly.
Our fourth assumption on the primitives of our model imposes that the proxy variables are exogenous.
Assumption H.2
$Cov\left(Z_{k},\epsilon\right)=0$ for any $k\in\left\{1,2\right\}$.
Fix $k\in\left\{1,2\right\}$ arbitrarily and note that
$Cov\left(U_{k},\epsilon_{res}\right)$ $=Cov\left(U_{k},\epsilon\right)$
$=Cov\left(Z_{k,res}-X_{res}^{*},\epsilon\right)$
$=Cov\left(Z_{k}-W\cdot\delta_{k}-X^{*}+W\cdot\delta_{X},\epsilon\right)$
$\hskip 20.0pt\text{where }\delta_{k}\coloneqq\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot Z_{k}\right]$
$\hskip 20.0pt\text{and }\delta_{X}\coloneqq\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot X^{*}\right]$
$=Cov\left(Z_{k}-X^{*}-W\cdot\left(\delta_{k}-\delta_{X}\right),\epsilon\right)$
$=Cov\left(Z_{k},\epsilon\right)-Cov\left(X^{*},\epsilon\right)-Cov\left(W,\epsilon\right)\cdot\left(\delta_{k}-\delta_{X}\right)$
$=0.$
Hence, Assumptions H.1 and H.2 implies Assumption 4.
Lastly, since $U_{1}$ and $U_{2}$ are defined based on model primitives, our fifth assumption is Assumption 5 directly.
Appendix I Sources of Measurement Error in FAO-GAEZ data
In this appendix, we discuss potential sources of measurement error in the FAO-GAEZ data. Our main goal is arguing the plausibility of Assumption 5.
Consider the following model describing how FAO-GAEZ version $v$ proxies for agricultural productivity of crop $c$
$Z_{v}=f_{v}(M+\eta_{v},\theta+\zeta_{v})$ (I.1)
where $f_{v}$ is a known function, $M$ denotes a vector of predictors (e.g., true precipitation and true temperature), and $\theta$ is a vector of parameters. Importantly, $M$ and $\theta$ are permanent objects in the sense that they do not change across versions. Since these permanent objects are not perfectly observed, we include two error terms that vary across versions: $\zeta_{v}$ captures calibration error and $\eta_{v}$ captures measurement error in the predictors.
Thus, the measurement error in version $v$ can be decomposed as:
$\begin{split}U_{v}&=f_{v}(M+\eta_{v},\theta+\zeta_{v})-X^{*}\\ &=\underbrace{f_{v}(M+\eta_{v},\theta+\zeta_{v})-f_{v}(M,\theta)}_{\text{imperfect data term}}+\underbrace{f_{v}(M,\theta)-X^{*}}_{\text{functional form misspecification}}\end{split}$ (I.2)
Intuitively, the first term $\left(f_{v}(M+\eta_{v},\theta+\zeta_{v})-f_{v}(M,\theta)\right)$ captures measurement error arising from imprecise input data. This term will be positively correlated across versions if the input data presents worse quality in developing regions over time. For example, developing regions may have a smaller number of weather stations and may rely more heavily on interpolation methods to estimate the predictors $M$, implying that $\eta_{v}$ is correlated across versions of the FAO-GAEZ project.
The second term $\left(f_{v}(M,\theta)-X^{*}\right)$ captures measurement error arising from functional form misspecification. This type of error occurs because researchers do not know the true model $f^{*}$ connecting predictors and parameters to crop productivity $X^{*}=f^{*}\left(M,\theta\right)$. This term will be positively correlated across versions if the proxy model $f_{v}$ is not updated heavily across versions of the FAO-GAEZ project. In particular, this term will be perfectly correlated across versions if $f_{v}$ is the same function for different values of $v$.
Lastly, this decomposition illustrates why Assumption 5 is plausible. A sufficient condition for $Cov(U_{1},U_{2})\geq 0$ is (i) persistence on both the imperfect data term and the misspecification term across versions and (ii) a non-negative covariance between these two terms across versions. According to the last two paragraphs, condition (i) is plausible in our setting. Furthermore, condition (ii) may arise if predictor measurement and functional-form assumptions are less accurate in developing regions. The latter is plausible in our context because scientific investment in developing regions is smaller.